### v2

# Xing Huang a), Feng-Yu Wang a)

Abstract
The log-Harnack inequality and Bismut formula are established for McKean-Vlasov SDEs with singularities in all (time, space, distribution) variables, where the drift satisfies a standard integrability condition in time-space, and may be discontinuous in the distance induced by any Dini function. The main results considerably improve the existing ones for the case where the drift is $L$-differentiable and Lipschitz continuous in distribution with respect to the quadratic Wasserstein distance.
AMS subject Classification:  60H10, 60B05.
Keywords: McKean-Vlasov SDEs, Log-Harnack inequality, Bismut formula, Dini function, Wasserstein distance.
1 Introduction
Let $\mathscr{P}$ be the set of all probability measures on $\mathbb{R}^{d}$ equipped with the weak topology, and let $W_{t}$ be an $m$-dimensional Brownian motion on a complete filtration probability space $(\Omega,\{\mathscr{F}_{t}\}_{t\geq 0},\mathscr{F},\mathbb{P})$. Consider the following McKean-Vlasov SDE on $\mathbb{R}^{d}$:
(1.1) $\text{\rm{d}}X_{t}=b_{t}(X_{t},\mathscr{L}_{X_{t}})\text{\rm{d}}t+\sigma_{t}(X_{t})\text{\rm{d}}W_{t},\ \ t\in[0,T],$
where $T>0$ is a fixed time, $\mathscr{L}_{X_{t}}$ is the distribution of $X_{t}$, and
$b:[0,T]\times\mathbb{R}^{d}\times\mathscr{P}\rightarrow\mathbb{R}^{d},\ \ \sigma:[0,T]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}\otimes\mathbb{R}^{m}$
are measurable. Because of its wide applications, this type SDE has been intensively investigated, see for instance [4, 5, 8, 9, 16, 17, 22] and the survey [11].
In this paper, we study the regularity of (1.1) for distributions in
$\mathscr{P}_{k}:=\big{\{}\mu\in\mathscr{P}:\|\mu\|_{k}:=\mu(|\cdot|^{k})^{\frac{1}{k}}<\infty\big{\}},\ \ k\in(1,\infty).$
Note that $\mathscr{P}_{k}$ is a Polish space under the Wasserstein distance
$\mathbb{W}_{k}(\mu,\nu)=\inf_{\pi\in\mathscr{C}(\mu,\nu)}\bigg{(}\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}|x-y|^{k}\pi(\text{\rm{d}}x,\text{\rm{d}}y)\bigg{)}^{\frac{1}{k}},$
where $\mathscr{C}(\mu,\nu)$ is the set of all couplings of $\mu$ and $\nu$. The SDE (1.1) is called well-posed for distributions in $\mathscr{P}_{k}$, if for any initial value $X_{0}$ with $\mathscr{L}_{X_{0}}\in\mathscr{P}_{k}$ (respectively, any initial distribution $\gamma\in\mathscr{P}_{k}$), it has a unique solution (respectively, a unique weak solution) $X=(X_{t})_{t\in[0,T]}$ such that $\mathscr{L}_{X_{\cdot}}:=(\mathscr{L}_{X_{t}})_{t\in[0,T]}\in C([0,T];\mathscr{P}_{k}).$ In this case, for any $\gamma\in\mathscr{P}_{k}$, let $P_{t}^{*}\gamma=\mathscr{L}_{X_{t}^{\gamma}}$ for the solution $X_{t}^{\gamma}$ with $\mathscr{L}_{X_{0}^{\gamma}}=\gamma$. We study the regularity of the map
$\mathscr{P}_{k}\ni\gamma\mapsto P_{t}f(\gamma):=\mathbb{E}[f(X_{t}^{\gamma})]=\int_{\mathbb{R}^{d}}f\text{\rm{d}}\{P_{t}^{*}\gamma\}$
for $t\in(0,T]$ and $f\in\mathscr{B}_{b}(\mathbb{R}^{d}),$ where $\mathscr{B}_{b}(\mathbb{R}^{d})$ is the space of bounded measurable functions on $\mathbb{R}^{d}$.
As powerful tools characterizing the regularity in distribution for stochastic systems, the dimension-free Harnack inequality due to [24], the log-Harnack inequality introduced in [25], and the Bismut (also called Bismut-Elworthy-Li) formula developed from [6, 10], have been intensively investigated. See for instance the monograph [26] for an account of related study on SPDEs.
In recent years, the log-Harnack inequality and Bismut type formula have also been established for McKean-Vlasov SDEs with coefficients regular in the distribution variable. Below we present a brief summary.
Write $b_{t}(x,\mu)=b_{t}^{(0)}(x)+b_{t}^{(1)}(x,\mu)$. According to [28], if $b^{(0)}$ satisfies some integrability condition on $(t,x)$, and there exists a constant $K_{b}\geq 0$ such that
$|b_{t}^{(1)}(x,\mu)-b_{t}^{(1)}(y,\nu)|\leq K_{b}(|x-y|^{2}+\mathbb{W}_{2}(\mu,\nu)^{2}),\ \ (x,\mu),(y,\nu)\in\mathbb{R}^{d}\times\mathscr{P}_{2},t\in[0,T],$
then there exists a constant $c>0$ such that the log-Harnack inequality
$P_{t}\log f(\tilde{\gamma})\leq\log P_{t}f(\gamma)+\frac{c}{t}\mathbb{W}_{2}(\gamma,\tilde{\gamma})^{2},\ \ t\in(0,T],\ f\in\mathscr{B}_{b}^{+}(\mathbb{R}^{d}),\ \gamma,\tilde{\gamma}\in\mathscr{P}_{2}$
holds, where $\mathscr{B}_{b}^{+}(\mathbb{R}^{d})$ is the space of positive elements in $\mathscr{B}_{b}(\mathbb{R}^{d})$. This inequality is equivalent to the entropy-cost inequality
${\rm Ent}(P_{t}^{*}\gamma|P_{t}^{*}\tilde{\gamma})\leq\frac{c}{t}\mathbb{W}_{2}(\gamma,\tilde{\gamma})^{2},\ \ t\in(0,T],\gamma,\tilde{\gamma}\in\mathscr{P}_{2},$
where ${\rm Ent}$ is the relative entropy, i.e. for any $\mu,\nu\in\mathscr{P}$, ${\rm Ent}(\nu|\mu):=\infty$ if $\nu$ is not absolutely continuous with respect to $\mu$, while
${\rm Ent}(\nu|\mu):=\mu(\rho\log\rho)=\int_{\mathbb{R}^{d}}(\rho\log\rho)\text{\rm{d}}\mu,\ \ \text{if}\ \rho:=\frac{\text{\rm{d}}\nu}{\text{\rm{d}}\mu}\ \text{exists. }$.
See also [14, 20, 27] for log-Harnack inequalities with more regular $b^{(0)}$, and see [18] for the dimension-free Harnack inequality with power.
If furthermore $b_{t}^{(1)}(x,\mu)$ is $L$-differentiable in $\mu\in\mathscr{P}_{k}$ (see Definition 3.1 below), the following Bismut type formula has been established in [29] for the intrinsic derivative $D^{I}_{\phi}$ (see Definition 3.1 below):
$D_{\phi}^{I}P_{t}f(\mu)=\mathbb{E}\big{[}f(X_{t}^{\mu})M_{t}^{\mu,\phi}\big{]},\ \ t\in(0,T],f\in\mathscr{B}_{b}(\mathbb{R}^{d}),\mu\in\mathscr{P}_{k},\phi\in L^{k}(\mathbb{R}^{d}\rightarrow\mathbb{R}^{d};\mu),$
where $M_{t}^{\mu,\phi}$ is an explicit martingale. See [3, 5, 12, 19] for earlier results with more regular $b^{(0)}$. See [2] for the case where $\mu=\delta_{x}$ is the Dirac measure at $x\in\mathbb{R}^{d}$, and see [21, 23] for a less explicit Bismut formula involving in the inverse of the Malliavin matrix of the solution.
We emphasize that existing results on log-Harnack inequality and Bismut formula for McKean-Vlasov SDEs only apply to the case with coefficients regular in the distribution variable, i.e. either $\mathbb{W}_{2}$-Lipschitz continuous or $L$-differentiable. The reason is that the Zvonkin transform technique [33] used in these references only kills singularities in the time-spatial variables $(t,x)$, but not the distribution variable.
On the other hand, a derivative estimate has been presented in [7] for the heat kernel when the drift is of type $b_{t}(x,\mu(V))$, where $V$ is a Hölder continuous function, and $\mu(V):=\int_{\mathbb{R}^{d}}V\text{\rm{d}}\mu$. In this case, the drift is only Lipschitz continuous in distribution with respect to
$\mathbb{W}_{\varepsilon}(\mu,\nu):=\sup\big{\{}\big{|}\mu(f)-\nu(f)\big{|}:|f(x)-f(y)|\leq|x-y|^{\varepsilon}\big{\}}$
rather than $\mathbb{W}_{1}$, and hence also has certain singularity in the distribution variable. This result encourages us to establish the log-Harnack inequality and Bismut formula for McKean-Vlasov SDEs with coefficients singular in all time-spatial-distribution variables.
Indeed, we will establish the log-Harnack inequality and Bismut formula for McKean-Vlasov SDEs with stronger singularity in the distribution variable: the drift is only Lipschitz continuous with respect to
$\mathbb{W}_{\alpha}(\mu,\nu):=\sup\big{\{}\big{|}\mu(f)-\nu(f)\big{|}:|f(x)-f(y)|\leq\alpha(|x-y|)\big{\}},$
where $\alpha$ is the square root of a Dini function, i.e. $\alpha$ is concave with $\alpha(0)=0,\alpha^{\prime}(s)>0$ for $s>0$, and $\int_{0}^{1}\frac{\alpha(s)^{2}}{s}\text{\rm{d}}s<\infty$ (Dini condition for $\alpha^{2}$). Thus, the continuity in the distribution variable is even weaker than the Dini continuity, so that the existing study in the literature is considerably improved.
To make this improvement, we develop a new technique by establishing a-priori derivative estimates and Bismut formula for the decoupled singular SDEs. The key point is to find a constant $c>0$ such that
(1.2) $\mathbb{W}_{\alpha}(P_{t}^{\ast}\gamma,P_{t}^{\ast}\tilde{\gamma})\leq c\mathbb{W}_{2}(\tilde{\gamma},\gamma)\frac{\alpha(t^{\frac{1}{2}})}{\sqrt{t}},\ \ \gamma,\tilde{\gamma}\in\mathscr{P}_{2},$
and for the drift of type $b_{t}(x,\mu(V))$ with $[V]_{\alpha}:=\sup_{x\neq y}\frac{|V(x)-V(y)|}{\alpha(|x-y|)}\leq 1$,
(1.3) $\|D^{I}P_{t}V(\mu)\|_{L^{\frac{k}{k-1}}(\mu)}\leq\frac{c\,\alpha(t^{\frac{1}{2}})}{\sqrt{t}},\ \ t\in(0,T],\mu\in\mathscr{P}_{k},$
where $D^{I}$ is the intrinsic derivative (see Definition 3.1 below).
2 Log-Harnack Inequality
Since $\mathbb{W}_{2}$ is involved in the log-Harnack inequality, in this section we mainly consider (1.1) for distributions in $\mathscr{P}_{2}$, but the drift may be discontinuous in $\mathbb{W}_{k}$ for any $k>0$. We first state the concrete assumption and the main result on the log-Harnack inequality, then present a complete proof in a separate subsection.
2.1 Assumption and main result
We will allow $b_{t}(x,\cdot)$ to be merely Lipschitz continuous in the sum of $\mathbb{W}_{2}$ and the Wasserstein distance induced by the square root of a Dini function.
Let $\alpha$ be in the following class
$\mathscr{A}:=\bigg{\{}\alpha:[0,\infty)\rightarrow[0,\infty)\ \text{is\ increasing\ and\ concave,\ }\alpha(0)=0,\ \int_{0}^{1}\frac{\alpha(r)^{2}}{r}\text{\rm{d}}r\in(0,\infty)\bigg{\}},$
where $\int_{0}^{1}\frac{\alpha(r)^{2}}{r}\text{\rm{d}}r<\infty$ is the Dini condition for $\alpha^{2}$. For a (real or Banach valued) function $f$, let
$[f]_{\alpha}:=\sup_{x\neq y}\frac{|f(x)-f(y)|}{\alpha(|x-y|)}$
be its continuity modulus in $\alpha$. Define the Wasserstein distance induced by $\alpha$:
$\mathbb{W}_{\alpha}(\mu,\nu):=\sup_{[f]_{\alpha}\leq 1}|\mu(f)-\nu(f)|,\ \ \mu,\nu\in\mathscr{P}_{\alpha}:=\big{\{}\mu\in\mathscr{P}:\mu(\alpha(|\cdot|))<\infty\big{\}},$
where $f$ are real functions and $\mu(f)=\int_{\mathbb{R}^{d}}f\text{\rm{d}}\mu$.
By the concavity of $\alpha$, $\mathbb{W}_{\alpha}$ is a complete distance on $\mathscr{P}_{\alpha}$, $\mathscr{P}_{k}\subset\mathscr{P}_{\alpha}$ for $k\geq 1$, and
(2.1) $\alpha(s+t)\leq\alpha(s)+\alpha(t),\ \ \alpha(rt)\leq r\alpha(t),\ \ \\ \ s,t>0,r\geq 1.$
These inequalities follow from $\alpha(0)=0$ and the decreasing monotonicity of $\alpha^{\prime}$ such that
$\alpha^{\prime}(s+t)\leq\alpha^{\prime}(s),\ \ \frac{\text{\rm{d}}}{\text{\rm{d}}t}\alpha(rt)=r\alpha^{\prime}(rt)\leq r\alpha^{\prime}(t),\ \ s,t\geq 0,r\geq 1.$
Since $\alpha$ is increasing with $\int_{0}^{1}\frac{\alpha(s)^{2}}{s}\text{\rm{d}}s>0,$ we have $\alpha(1)>0$ so that the second estimate in (2.1) with $r=t^{-1}$ yields
(2.2) $\alpha(t)\geq\alpha(1)t>0,\ \ \ t\in(0,1].$
To measure the singularity in $(t,x)\in[0,T]\times\mathbb{R}^{d}$, we recall locally integrable functional spaces presented in [30]. For any $t>s\geq 0$ and $p,q\in(1,\infty]$, we write $f\in\tilde{L}_{p}^{q}([s,t])$ if $f:[s,t]\times\mathbb{R}^{d}\rightarrow\mathbb{R}$ is measurable with
$\|f\|_{\tilde{L}_{p}^{q}([s,t])}:=\sup_{y\in\mathbb{R}^{d}}\bigg{\{}\int_{s}^{t}\bigg{(}\int_{B(y,1)}|f(r,x)|^{p}\text{\rm{d}}x\bigg{)}^{\frac{q}{p}}\text{\rm{d}}r\bigg{\}}^{\frac{1}{q}}<\infty,$
where $B(y,1):=\{x\in\mathbb{R}^{d}:|x-y|\leq 1\}$ is the unit ball centered at the point $y$. When $s=0$, we simply denote
$\tilde{L}_{p}^{q}(t)=\tilde{L}_{p}^{q}([0,t]),\ \ \|f\|_{\tilde{L}_{p}^{q}(t)}=\|f\|_{\tilde{L}_{p}^{q}([0,t])}.$
We take $(p,q)$ from the space
$\mathscr{K}:=\Big{\{}(p,q)\in(2,\infty]^{2},\ \frac{d}{p}+\frac{2}{q}<1\Big{\}},$
and make the following assumption where $\nabla$ is the gradient in $x\in\mathbb{R}^{d}$.
  1. (A)
    There exist $\alpha\in\mathscr{A}$, $k\in(1,\infty),K\in(0,\infty),l\in\mathbb{N},$ and
[…]
    such that the following conditions hold.
  2. $(A_{1})$
    $(\sigma_{t}\sigma^{*}_{t})(x)$ is invertible and $\sigma_{t}(x)$ is weakly differentiable in $x$ such that
    $\|\sigma\sigma^{*}\|_{\infty}+\|(\sigma\sigma^{*})^{-1}\|_{\infty}<\infty,\ \ |\nabla\sigma|\leq\sum_{i=1}^{l}f_{i},$
    $\lim_{\varepsilon\downarrow 0}\sup_{t\in[0,T],|x-x^{\prime}|\leq\varepsilon}\|(\sigma_{t}\sigma_{t}^{\ast})(x)-(\sigma_{t}\sigma_{t}^{\ast})(x^{\prime})\|=0.$
  3. $(A_{2})$
    $b_{t}(x,\mu)=b_{t}^{(0)}(x)+b_{t}^{(1)}(x,\mu),$ where for any $t\in[0,T],x,y\in\mathbb{R}^{d},\mu,\nu\in\mathscr{P}_{k},$
    $|b^{(0)}_{t}(x)|\leq f_{0}(t,x),\ \ |b_{t}^{(1)}(0,\delta_{0})|\leq K,$
    $|b_{t}^{(1)}(x,\mu)-b_{t}^{(1)}(y,\nu)|\leq K\big{\{}|x-y|+\mathbb{W}_{\alpha}(\mu,\nu)+\mathbb{W}_{k}(\mu,\nu)\big{\}}.$
We first observe that (A) implies the well-posedness of (1.1) for distributions in $\mathscr{P}_{k}$. Since $\alpha(0)=0$ and $\alpha$ is concave, there exists a constant $c>0$ such that
$\sup_{[f]_{\alpha}\leq 1}|f(x)-f(0)|\leq\alpha(|x|)\leq c+c|x|^{k},\ \ x\in\mathbb{R}^{d}.$
Thus,
(2.3) $\frac{1}{c}\mathbb{W}_{\alpha}(\mu,\nu)\leq\mathbb{W}_{k,var}(\mu,\nu):=\sup_{|f|\leq 1+|\cdot|^{k}}|\mu(f)-\nu(f)|.$
So, by [28, Theorem 3.1(1)] for $D=\mathbb{R}^{d}$, under assumption (A), (1.1) is well-posed for distributions in $\mathscr{P}_{k}$, and for any $n\geq 1$ there exists a constant $c_{n}>0$ such that
$\mathbb{E}\Big{[}\sup_{t\in[0,T]}|X_{t}|^{n}\Big{|}\mathscr{F}_{0}\Big{]}\leq c_{n}(1+|X_{0}|^{n}).$
Consequently,
(2.4) $\sup_{t\in[0,T]}\|P_{t}^{*}\gamma\|_{n}^{n}\leq\mathbb{E}\Big{[}\sup_{t\in[0,T]}|X_{t}^{\gamma}|^{n}\Big{]}\leq c_{n}(1+\|\gamma\|_{n}^{n}),\ \ n\geq 1.$
Theorem 2.1.
Assume (A) with $k=2$. Then there exists a constant $c>0$ such that
(2.5) ${\rm Ent}(P_{t}^{\ast}\gamma|P_{t}^{\ast}\tilde{\gamma})\leq\frac{c}{t}\,\mathbb{W}_{2}(\gamma,\tilde{\gamma})^{2},\ \ t\in(0,T],\gamma,\tilde{\gamma}\in\mathscr{P}_{2}.$
Example 2.2.
Let $h:\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}$ satisfy
$|h(x_{1},y_{1})-h(x_{2},y_{2})|\leq K_{h}|x_{1}-x_{2}|+\alpha(|y_{1}-y_{2}|),\ \ x_{1},x_{2},y_{1},y_{2}\in\mathbb{R}^{d}$
for some $\alpha\in\mathscr{A}$ and $K_{h}\geq 0$. Then $b_{t}^{(1)}(x,\mu)=\int_{\mathbb{R}^{d}}h(x,y)\mu(\text{\rm{d}}y)$ satisfies ($A_{2}$).
2.2 Proof of Theorem 2.1
Although in Theorem 2.1 we assume (A) for $k=2$, for later use we will also consider general $k\in(1,\infty)$. For any $\gamma\in\mathscr{P}_{k}$, consider the decoupled SDE of (1.1):
(2.6) $\text{\rm{d}}X^{x,\gamma}_{t}=b_{t}(X^{x,\gamma}_{t},P_{t}^{*}\gamma)\text{\rm{d}}t+\sigma_{t}(X^{x,\gamma}_{t})\text{\rm{d}}W_{t},\ \ X^{x,\gamma}_{0}=x.$
To characterize properties of (2.6), consider the following PDE for $u:[0,T]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ :
(2.7) $\frac{\partial}{\partial t}u_{t}(x)+(\mathscr{L}_{t}^{\gamma}u_{t})(x)+b_{t}^{(0)}(x)=\lambda u_{t}(x),\ \ u_{T}=0,$
where $\lambda>0$ is a constant, and
(2.8) $\mathscr{L}_{t}^{\gamma}:=\frac{1}{2}\mathrm{tr}\big{\{}(\sigma_{t}\sigma_{t}^{\ast})\nabla^{2}\big{\}}+\nabla_{b_{t}(\cdot,P_{t}^{*}\gamma)}.$
By [32, Theorem 2.1] and (A), for large enough constants $\lambda,c>0$ independent of $\gamma$, (2.7) has a unique solution ${u}^{\lambda,\gamma}$ satisfying
(2.9) $\|u^{\lambda,\gamma}\|_{\infty}+\|\nabla{u}^{\lambda,\gamma}\|_{\infty}\leq\frac{1}{5},\ \ \|\nabla^{2}{u}^{\lambda,\gamma}\|_{\tilde{L}_{p_{0}}^{q_{0}}(T)}\leq c.$
So, for any $t\in[0,T]$,
(2.10) $x\mapsto\Theta^{\lambda,\gamma}_{t}(x):=x+{u}^{\lambda,\gamma}_{t}(x),\ \ x\in\mathbb{R}^{d}$
is a homeomorphism on $\mathbb{R}^{d}$.
Moreover, for any $\gamma\in\mathscr{P}_{k}$, consider
(2.11) $\text{\rm{d}}\theta^{\lambda,\gamma}_{t}(x)=b^{(1)}_{t}((\Theta^{\lambda,\gamma}_{t})^{-1}(\theta^{\lambda,\gamma}_{t}(x)),P_{t}^{\ast}\gamma)\text{\rm{d}}t,\ \ t\in[0,T],\theta^{\lambda,\gamma}_{0}(x)=\Theta^{\lambda,\gamma}_{0}(x),x\in\mathbb{R}^{d},$
and let
(2.12) $\tilde{\theta}^{\lambda,\gamma}_{t}(x)=(\Theta^{\lambda,\gamma}_{t})^{-1}(\theta^{\lambda,\gamma}_{t}(x)),\ \ t\in[0,T],x\in\mathbb{R}^{d}.$
Then we have
(2.13) $\text{\rm{d}}\Theta^{\lambda,\gamma}_{t}(\tilde{\theta}^{\lambda,\gamma}_{t}(x))=b^{(1)}_{t}(\tilde{\theta}^{\lambda,\gamma}_{t}(x),P_{t}^{\ast}\gamma)\text{\rm{d}}t,\ \ t\in[0,T],\tilde{\theta}^{\lambda,\gamma}_{0}(x)=x\in\mathbb{R}^{d}.$
Lemma 2.3.
Let $\sigma$ and $b$ satisfy (A). The the following assertions hold.
(1) The SDE (2.6) is well-posed and for any $v\in\mathbb{R}^{d},\gamma\in\mathscr{P}_{k},x\in\mathbb{R}^{d}$, the limit
$\nabla_{v}X_{t}^{x,\gamma}:=\lim_{\varepsilon\downarrow 0}\frac{X_{t}^{x+\varepsilon v,\gamma}-X_{t}^{x,\gamma}}{\varepsilon},\ \ \ \ t\in[0,T]$
exists in $L^{p}(\Omega\rightarrow C([0,T];\mathbb{R}^{d});\mathbb{P})$ for any $p\geq 1$, and there exists a constant $c_{p}>0$ such that
(2.14) $\mathbb{E}\Big{[}\sup_{t\in[0,T]}|\nabla_{v}X_{t}^{x,\gamma}|^{p}\Big{]}\leq c_{p}|v|^{p},\ \ v\in\mathbb{R}^{d},\gamma\in\mathscr{P}_{k},x\in\mathbb{R}^{d}.$
Moreover, letting $\zeta_{s}:=\sigma_{s}^{*}(\sigma_{s}\sigma_{s}^{*})^{-1}$ and $P_{t}^{\gamma}$ be the associated Markov semigroup, i.e.
$P_{t}^{\gamma}f(x):=\mathbb{E}[f(X_{t}^{x,\gamma})],\ \ t\in[0,T],x\in\mathbb{R}^{d},f\in\mathscr{B}_{b}(\mathbb{R}^{d}),$
then the Bismut formula for $P^{\gamma}_{t}$ holds:
$\nabla_{v}P^{\gamma}_{t}f(x)=\mathbb{E}\left[f(X^{x,\gamma}_{t})\int_{0}^{t}\frac{1}{t}\big{\langle}\zeta_{s}(X^{x,\gamma}_{s})\nabla_{v}X^{x,\gamma}_{s},\text{\rm{d}}W_{s}\big{\rangle}\right],\ \ x,v\in\mathbb{R}^{d},\gamma\in\mathscr{P}_{k},t\in(0,T].$
(2) For any $p\geq 1$, there exists a constant $c_{p}>0$ such that
(2.15) $\mathbb{E}[|X^{x,\gamma}_{t}-\tilde{\theta}^{\lambda,\gamma}_{t}(x)|^{p}]\leq c_{p}t^{\frac{p}{2}},\ \ t\in[0,T],x\in\mathbb{R}^{d},\gamma\in\mathscr{P}_{k},$
Proof.
By [29, Theorem 2.1] for $\beta_{s}=\frac{s}{t}$, (A) implies (1). It remains to prove (2.15). To this end, we use Zvonkin’s transform.
By Itô’s formula (see [32, Lemma 3.3]), we derive
(2.16) $\begin{split}\text{\rm{d}}\Theta^{\lambda,\gamma}_{t}(X^{x,\gamma}_{t})&=\big{\{}\lambda u^{\lambda,\gamma}_{t}(X^{x,\gamma}_{t})+b_{t}^{(1)}(X^{x,\gamma}_{t},P_{t}^{*}\gamma)\big{\}}\text{\rm{d}}t+\big{\{}(\nabla\Theta_{t}^{\lambda,\gamma})\sigma_{t}\big{\}}(X^{x,\gamma}_{t})\,\text{\rm{d}}W_{t}.\end{split}$
By (A), (2.9), there exists a constant $C>1$ such that
$C^{-1}|X^{x,\gamma}_{t}-\tilde{\theta}^{\lambda,\gamma}_{t}(x)|\leq\big{|}\Theta^{\lambda,\gamma}_{t}(X^{x,\gamma}_{t})-\Theta^{\lambda,\gamma}_{t}(\tilde{\theta}^{\lambda,\gamma}_{t}(x))\big{|}\leq C|X^{x,\gamma}_{t}-\tilde{\theta}^{\lambda,\gamma}_{t}(x)|,$
$\big{|}b_{t}^{(1)}(X^{x,\gamma}_{t})-b_{t}^{(1)}(\tilde{\theta}^{\lambda,\gamma}_{t}(x))\big{|}\leq C|X^{x,\gamma}_{t}-\tilde{\theta}^{\lambda,\gamma}_{t}(x)|,$
$\big{|}\lambda u^{\lambda,\gamma}_{t}(X^{x,\gamma}_{t})\big{|}+\big{\|}\big{\{}(\nabla\Theta_{t}^{\lambda,\gamma})\sigma_{t}\big{\}}(X^{x,\gamma}_{t})\big{\|}\leq C,\ \ (t,x,\gamma)\in[0,T]\times\mathbb{R}^{d}\times\mathscr{P}_{k}.$
This together with (2.13), (2.16) and Gronwall’s inequality implies (2.15). ∎
To verify (2.5), in the following Lemma 2.4 and Lemma 2.6 we will prove
(2.17) $\int_{0}^{t}\{\mathbb{W}_{\alpha}(P_{s}^{*}\gamma,P_{s}^{*}\tilde{\gamma})+\mathbb{W}_{2}(P_{s}^{*}\gamma,P_{s}^{*}\tilde{\gamma})\}^{2}\text{\rm{d}}s\leq c\mathbb{W}_{2}(\gamma,\tilde{\gamma})^{2},\ \ t\in[0,T],\gamma,\tilde{\gamma}\in\mathscr{P}_{2}$
for some constant $c>0$.
Lemma 2.4.
Assume (A). Then there exists a constant $c>0$ such that
(2.18) $\mathbb{W}_{k}(P_{t}^{\ast}\gamma,P_{t}^{\ast}\tilde{\gamma})\leq c\mathbb{W}_{k}(\gamma,\tilde{\gamma})+c\int_{0}^{t}\mathbb{W}_{\alpha}(P_{s}^{\ast}\gamma,P_{s}^{\ast}\tilde{\gamma})\text{\rm{d}}s,\ \ t\in[0,T],\gamma,\tilde{\gamma}\in\mathscr{P}_{k}.$
Proof.
We take $\mathscr{F}_{0}$-measurable random variables $X_{0}^{\gamma},X_{0}^{\tilde{\gamma}}$ such that
(2.19) $\mathscr{L}_{X_{0}^{\gamma}}=\gamma,\ \ \mathscr{L}_{X_{0}^{\tilde{\gamma}}}=\tilde{\gamma},\ \ \mathbb{W}_{k}(\gamma,\tilde{\gamma})^{k}=\mathbb{E}[|X_{0}^{\gamma}-X_{0}^{\tilde{\gamma}}|^{k}].$
Recall that $\Theta^{\lambda,\gamma}_{t}$ is defined in (2.10). By (2.7), (2.8) and Itô’s formula, we derive
(2.20) $\begin{split}\text{\rm{d}}\Theta^{\lambda,\gamma}_{t}(X^{\gamma}_{t})&=\big{\{}\lambda u^{\lambda,\gamma}_{t}(X^{\gamma}_{t})+b_{t}^{(1)}(X^{\gamma}_{t},P_{t}^{*}\gamma)\big{\}}\text{\rm{d}}t+\big{\{}(\nabla\Theta_{t}^{\lambda,\gamma})\sigma_{t}\big{\}}(X^{\gamma}_{t})\,\text{\rm{d}}W_{t},\end{split}$
and
(2.21) $\begin{split}&\text{\rm{d}}\Theta^{\lambda,\gamma}_{t}(X^{\tilde{\gamma}}_{t})=\big{\{}\lambda u^{\lambda,\gamma}_{t}(X^{\tilde{\gamma}}_{t})+b_{t}^{(1)}(X^{\tilde{\gamma}}_{t},P_{t}^{*}\gamma)\big{\}}\text{\rm{d}}t\\ &\qquad+\nabla_{b_{t}(X^{\tilde{\gamma}}_{t},P_{t}^{\ast}\tilde{\gamma})-b_{t}(X^{\tilde{\gamma}}_{t},P_{t}^{\ast}\gamma)}\Theta^{\lambda,\gamma}_{t}(X^{\tilde{\gamma}}_{t})\text{\rm{d}}t+\big{\{}(\nabla\Theta_{t}^{\lambda,\gamma})\sigma_{t}\big{\}}(X^{\tilde{\gamma}}_{t})\,\text{\rm{d}}W_{t}.\end{split}$
Combining this with (2.20) and (A), we prove the desired estimate by using the maximal functional inequality, Khasminskii’s estimate and stochastic Gronwall’s inequality, see for instance the proof of [15, Lemma 2.1] for details. Below we simply outline the procedure.
By $(A_{2})$ we have
$|b_{t}(X^{\tilde{\gamma}}_{t},P_{t}^{\ast}\tilde{\gamma})-b_{t}(X^{\tilde{\gamma}}_{t},P_{t}^{\ast}\gamma)|+|b_{t}^{(1)}(X^{\tilde{\gamma}}_{t},P_{t}^{\ast}\gamma)-b_{t}^{(1)}(X^{\gamma}_{t},P_{t}^{\ast}\gamma)|$
$\leq K\big{\{}|X^{\gamma}_{t}-X^{\tilde{\gamma}}_{t}|+\mathbb{W}_{\alpha}(P_{t}^{\ast}\gamma,P_{t}^{*}\tilde{\gamma})+\mathbb{W}_{k}(P_{t}^{*}\gamma,P_{t}^{*}\tilde{\gamma})\big{\}}.$
Combining this with (2.20), (2.21), $(A_{1})$, the maximal functional inequality and Khasminskii’s estimate (see [30, Lemma 2.1 and Lemma 4.1]), we derive
$\text{\rm{d}}\big{|}\Theta^{\lambda,\gamma}_{t}(X^{\gamma}_{t})-\Theta^{\lambda,\gamma}_{t}(X^{\tilde{\gamma}}_{t})\big{|}^{k+1}\leq\text{\rm{d}}M_{t}+|X^{\gamma}_{t}-X^{\tilde{\gamma}}_{t}|^{k+1}\text{\rm{d}}\mathscr{L}_{t}$
$\qquad+c_{1}\big{\{}\mathbb{W}_{\alpha}(P_{t}^{\ast}\gamma,P_{t}^{*}\tilde{\gamma})+\mathbb{W}_{k}(P_{t}^{*}\gamma,P_{t}^{*}\tilde{\gamma})\big{\}}|\Theta^{\lambda,\gamma}_{t}(X^{\gamma}_{t})-\Theta^{\lambda,\gamma}_{t}(X^{\tilde{\gamma}}_{t})\big{|}^{k}\text{\rm{d}}t,$
where $c_{1}>0$ is a constant, $\mathscr{L}_{t}$ is an adapted increasing process with $\mathbb{E}[\text{\rm{e}}^{\delta\mathscr{L}_{T}}]<\infty$ for any $\delta>0$, and $M_{t}$ is a local martingale. Since (2.9) implies
$\frac{1}{2}|X^{\tilde{\gamma}}_{t}-X^{\gamma}_{t}|\leq|\Theta^{\lambda,\gamma}_{t}(X^{\gamma}_{t})-\Theta^{\lambda,\gamma}_{t}(X^{\tilde{\gamma}}_{t})\big{|}\leq 2|X^{\tilde{\gamma}}_{t}-X^{\gamma}_{t}|,$
by the stochastic Gronwall inequality (see [31, Lemma 3.7]), we find a constant $c_{2}>1$ such that
$\bigg{\{}\mathbb{E}\Big{[}\sup_{s\in[0,t]}|X^{\tilde{\gamma}}_{s}-X^{\gamma}_{s}|^{k}\Big{|}\mathscr{F}_{0}\Big{]}\bigg{\}}^{1+k^{-1}}-c_{2}|X_{0}^{\gamma}-X_{0}^{\tilde{\gamma}}|^{k+1}$
$\leq c_{2}\int_{0}^{t}\big{\{}\mathbb{W}_{\alpha}(P_{s}^{\ast}\gamma,P_{s}^{*}\tilde{\gamma})+\mathbb{W}_{k}(P_{s}^{*}\gamma,P_{s}^{*}\tilde{\gamma})\big{\}}\mathbb{E}\Big{[}|X^{\tilde{\gamma}}_{s}-X^{\gamma}_{s}|^{k}\Big{|}\mathscr{F}_{0}\Big{]}\text{\rm{d}}s,\ \ t\in[0,T].$
So, there exists a constant $c_{3}>0$ such that for any $t\in[0,T]$,
$\mathbb{E}\Big{[}\sup_{s\in[0,t]}|X^{\tilde{\gamma}}_{s}-X^{\gamma}_{s}|^{k}\Big{|}\mathscr{F}_{0}\Big{]}-c_{2}|X_{0}^{\gamma}-X_{0}^{\tilde{\gamma}}|^{k}$
$\leq c_{2}\bigg{(}\int_{0}^{t}\big{\{}\mathbb{W}_{\alpha}(P_{s}^{\ast}\gamma,P_{s}^{*}\tilde{\gamma})+\mathbb{W}_{k}(P_{s}^{*}\gamma,P_{s}^{*}\tilde{\gamma})\big{\}}\mathbb{E}\Big{[}|X^{\tilde{\gamma}}_{s}-X^{\gamma}_{s}|^{k}\Big{|}\mathscr{F}_{0}\Big{]}\text{\rm{d}}s\bigg{)}^{\frac{k}{k+1}}$
$\leq\frac{1}{2}\mathbb{E}\Big{[}\sup_{s\in[0,t]}|X^{\tilde{\gamma}}_{s}-X^{\gamma}_{s}|^{k}\Big{|}\mathscr{F}_{0}\Big{]}+c_{3}\bigg{(}\int_{0}^{t}\big{\{}\mathbb{W}_{\alpha}(P_{s}^{\ast}\gamma,P_{s}^{*}\tilde{\gamma})+\mathbb{W}_{k}(P_{s}^{*}\gamma,P_{s}^{*}\tilde{\gamma})\big{\}}\text{\rm{d}}s\bigg{)}^{k}.$
This together with (2.19) yields
$\mathbb{W}_{k}(P_{t}^{*}\gamma,P_{t}^{*}\tilde{\gamma})\leq\sup_{s\in[0,t]}\big{(}\mathbb{E}[|X^{\tilde{\gamma}}_{s}-X^{\gamma}_{s}|^{k}]\big{)}^{\frac{1}{k}}$
$\leq(2c_{2})^{\frac{1}{k}}\mathbb{W}_{k}(\gamma,\tilde{\gamma})+(2c_{3})^{\frac{1}{k}}\int_{0}^{t}\big{\{}\mathbb{W}_{\alpha}(P_{s}^{\ast}\gamma,P_{s}^{*}\tilde{\gamma})+\mathbb{W}_{k}(P_{s}^{*}\gamma,P_{s}^{*}\tilde{\gamma})\big{\}}\text{\rm{d}}s,\ \ t\in[0,T].$
By Gronwall’s inequality, this implies the desired estimate for some constant $c>0$. ∎
To build the estimate (1.2), we present the following Hölder inequality for concave functions.
Lemma 2.5.
Let $\alpha:[0,\infty)\rightarrow[0,\infty)$ be concave. For any non-negative random variables $\xi$ and $\eta$,
$\mathbb{E}[\alpha(\xi)\eta]\leq\|\eta\|_{L^{p}(\mathbb{P})}\alpha\Big{(}\|\xi\|_{L^{\frac{p}{p-1}}(\mathbb{P})}\Big{)},\ \ p>1.$
Consequently, for any random variable $\xi$ on $\mathbb{R}^{d}$, $f\in C(\mathbb{R}^{d};\mathbb{B})$ for a Banach space $(\mathbb{B},\|\cdot\|_{\mathbb{B}})$ with $[f]_{\alpha}<\infty$, and any real random variable $\eta$ with $\mathbb{E}[\eta]=0$,
(2.22) $\big{\|}\mathbb{E}[f(\xi)\eta]\big{\|}_{\mathbb{B}}\leq[f]_{\alpha}\|\eta\|_{L^{p}(\mathbb{P})}\alpha\Big{(}\|\xi-y\|_{L^{\frac{p}{p-1}}(\mathbb{P})}\Big{)},\ \ p>1,y\in\mathbb{R}^{d}.$
Proof.
Since the assertion holds trivially for $p=\infty$, we only prove for $p<\infty$. It suffices to prove for $\mathbb{E}[|\eta|^{p}]\in(0,\infty)$. Let $\mathbb{Q}:=\frac{\eta}{\mathbb{E}[\eta]}\mathbb{P}.$ By Jensen’s and Hölder’s inequalities, and using (2.1), we obtain
$\mathbb{E}[\alpha(\xi)\eta]=\mathbb{E}[|\eta|]\mathbb{E}_{\mathbb{Q}}[\alpha(\xi)]\leq\mathbb{E}[|\eta|]\alpha(\mathbb{E}_{\mathbb{Q}}[\xi])\leq\mathbb{E}[|\eta|]\alpha\bigg{(}\frac{(\mathbb{E}[|\eta|^{p}])^{\frac{1}{p}}}{\mathbb{E}[|\eta|]}\big{(}\mathbb{E}[|\xi|^{\frac{p}{p-1}}]\big{)}^{\frac{p-1}{p}}\bigg{)}$
$\leq\mathbb{E}[|\eta|]\bigg{\{}\frac{(\mathbb{E}[|\eta|^{p}])^{\frac{1}{p}}}{\mathbb{E}[|\eta|]}\ \alpha\Big{(}\big{(}\mathbb{E}[|\xi|^{\frac{p}{p-1}}]\big{)}^{\frac{p-1}{p}}\Big{)}\bigg{\}}=\big{(}\mathbb{E}[|\eta|^{p}]\big{)}^{\frac{1}{p}}\alpha\Big{(}\big{(}\mathbb{E}[|\xi|^{\frac{p}{p-1}}]\big{)}^{\frac{p-1}{p}}\Big{)}.$
Then the second inequality follows by noting that $\mathbb{E}[\eta]=0$ implies
$\big{\|}\mathbb{E}[f(\xi)\eta]\big{\|}_{\mathbb{B}}=\big{\|}\mathbb{E}[\{f(\xi)-f(y)\}\eta]\big{\|}_{\mathbb{B}}\leq[f]_{\alpha}\mathbb{E}[\alpha(|\xi-y|)|\eta|].$
Noting that $X_{t}^{x,\gamma}$ solves (1.1) if the initial value $x$ is random with distribution $\gamma$, by the standard Markov property of $X_{t}^{x,\gamma}$, we have
(2.23) $P_{t}f(\gamma):=\int_{\mathbb{R}^{d}}f(x)(P_{t}^{*}\gamma)(\text{\rm{d}}x)=\int_{\mathbb{R}^{d}}P^{\gamma}_{t}f(x)\gamma(\text{\rm{d}}x),\ \ f\in\mathscr{B}_{b}(\mathbb{R}^{d}).$
The following lemma provides an extension to (1.2), which combined with Lemma 2.4 implies the desired (2.17).
Lemma 2.6.
Assume (A). Then there exists a constant $c>0$ such that
$\mathbb{W}_{\alpha}(P_{t}^{\ast}\gamma,P_{t}^{\ast}\tilde{\gamma})\leq c\mathbb{W}_{k}(\tilde{\gamma},\gamma)\frac{\alpha(t^{\frac{1}{2}})}{\sqrt{t}},\ \ t\in(0,T],\gamma,\tilde{\gamma}\in\mathscr{P}_{k}.$
Consequently, there exists a constant $c>0$ such that for any $\gamma,\tilde{\gamma}\in\mathscr{P}_{k},$
(2.24) $\sup_{t\in[0,T]}\mathbb{W}_{k}(P_{t}^{\ast}\gamma,P_{t}^{\ast}\tilde{\gamma})\leq c\mathbb{W}_{k}(\gamma,\tilde{\gamma}).$
Proof.
It suffices to prove the first estimate, since it implies (2.24) according to Lemma 2.4.
Let $X_{0}^{\gamma}$ and $X_{0}^{\tilde{\gamma}}$ be in (2.19). For any $\varepsilon\in[0,2]$, let
$X_{0}^{\gamma^{\varepsilon}}:=X_{0}^{\gamma}+\varepsilon(X_{0}^{\tilde{\gamma}}-X_{0}^{\gamma}),\ \ \ \gamma^{\varepsilon}:=\mathscr{L}_{X_{0}^{\gamma^{\varepsilon}}},$
and let $X_{t}^{\gamma^{\varepsilon}}$ solve (1.1) with initial value $X_{0}^{\gamma^{\varepsilon}}.$ Then
(2.25) $\gamma^{\varepsilon}(|\cdot|)\leq 2\|\gamma\|_{k}+2\|\tilde{\gamma}\|_{k},\ \ \varepsilon\in[0,2],$
(2.26) $\mathbb{W}_{k}(\gamma^{\varepsilon},\gamma^{\varepsilon+r})^{k}\leq\mathbb{E}\big{[}|X_{0}^{\gamma^{\varepsilon}}-X_{0}^{\gamma^{\varepsilon+r}}|^{k}\big{]}=r^{k}\mathbb{W}_{k}(\gamma,\tilde{\gamma})^{k},\ \ \varepsilon,r\in[0,1].$
For any $\varepsilon\geq 0$, consider the SDE
(2.27) $\text{\rm{d}}X_{t}^{x,\gamma^{\varepsilon}}=b_{t}(X_{t}^{x,\gamma^{\varepsilon}},P_{t}^{*}\gamma^{\varepsilon})\text{\rm{d}}t+\sigma_{t}(X_{t}^{x,\gamma^{\varepsilon}})\text{\rm{d}}W_{t},\ \ X_{0}^{x,\gamma^{\varepsilon}}=x,t\in[0,T].$
For any $r\in(0,1)$, let
$\eta_{t}^{\varepsilon,r}=[\sigma_{t}^{\ast}(\sigma_{t}\sigma_{t}^{\ast})^{-1}](X^{x,\gamma^{\varepsilon}}_{t})[b_{t}(X^{x,\gamma^{\varepsilon}}_{t},P_{t}^{*}\gamma^{\varepsilon+r})-b_{t}(X^{x,\gamma^{\varepsilon}}_{t},P_{t}^{*}\gamma^{\varepsilon})],\ \ t\in[0,T].$
By (A), there exists a constant $c_{1}>0$ such that
(2.28) $\sup_{t\in[0,T]}|\eta^{\varepsilon,r}_{t}|\leq c_{1}\big{\{}\mathbb{W}_{\alpha}(P_{t}^{\ast}\gamma^{\varepsilon},P_{t}^{\ast}\gamma^{\varepsilon+r})+\mathbb{W}_{k}(P_{t}^{\ast}\gamma^{\varepsilon},P_{t}^{\ast}\gamma^{\varepsilon+r})\big{\}},\ \ r,\varepsilon\in[0,1].$
By Girsanov’s theorem,
$R^{\varepsilon,r}_{t}:=\exp\left\{\int_{0}^{t}\langle\eta^{\varepsilon,r}_{s},\text{\rm{d}}W_{s}\rangle-\frac{1}{2}\int_{0}^{t}|\eta^{\varepsilon,r}_{s}|^{2}\text{\rm{d}}s\right\},\ \ t\in[0,T]$
is a martingale, and
$W_{t}^{\varepsilon,r}=W_{t}-\int_{0}^{t}\eta_{s}^{\varepsilon,r}\text{\rm{d}}s,\ \ \ t\in[0,T]$
is a Brownian motion under the probability measure $\mathbb{Q}^{\varepsilon,r}:=R^{\varepsilon,r}_{T}\mathbb{P}.$ Rewrite (2.27) as
$\text{\rm{d}}X^{x,\gamma^{\varepsilon}}_{t}=b_{t}(X^{x,\gamma^{\varepsilon}}_{t},P_{t}^{*}\gamma^{\varepsilon+r})\text{\rm{d}}t+\sigma_{t}(X^{x,\gamma^{\varepsilon}}_{t})\text{\rm{d}}W^{\varepsilon,r}_{t},\ \ X^{x,\gamma^{\varepsilon}}_{0}=x,\ \ t\in[0,T].$
By the weak uniqueness we obtain
$\mathscr{L}_{\{X^{x,\gamma^{\varepsilon}}_{t}\}_{t\in[0,T]}|\mathbb{Q}^{\varepsilon,r}}=\mathscr{L}_{\{X^{x,\gamma^{\varepsilon+r}}_{t}\}_{t\in[0,T]}},$
where $\mathscr{L}_{\cdot|\mathbb{Q}^{\varepsilon,r}}$ is the law under $\mathbb{Q}^{\varepsilon,r}$, so that
$P_{t}^{\gamma^{\varepsilon+r}}f(x)-P^{\gamma^{\varepsilon}}_{t}f(x)=\mathbb{E}\left[f(X^{x,\gamma^{\varepsilon}}_{t})(R_{t}^{\varepsilon,r}-1)\right],\ \ f\in\mathscr{B}_{b}(\mathbb{R}^{d}),\varepsilon,r\in(0,1].$
Hence, by (2.23), we have
$P_{t}f(\gamma^{\varepsilon+r})-P_{t}f(\gamma^{\varepsilon})=\gamma^{\varepsilon+r}(P_{t}^{\gamma^{\varepsilon+r}}f)-\gamma^{\varepsilon}(P^{\gamma^{\varepsilon}}_{t}f)$
$=\gamma^{\varepsilon+r}(P_{t}^{\gamma^{\varepsilon+r}}f-P^{\gamma^{\varepsilon}}_{t}f)+\gamma^{\varepsilon+r}(P_{t}^{\gamma^{\varepsilon}}f)-\gamma^{\varepsilon}(P^{\gamma^{\varepsilon}}_{t}f)$
$=\int_{\mathbb{R}^{d}}\mathbb{E}\left[f(X^{x,\gamma^{\varepsilon}}_{t})(R_{t}^{\varepsilon,r}-1)\right]\gamma^{\varepsilon+r}(\text{\rm{d}}x)+\mathbb{E}\big{[}P_{t}^{\gamma^{\varepsilon}}f(X_{0}^{\gamma^{\varepsilon+r}})-P_{t}^{\gamma^{\varepsilon}}f(X_{0}^{\gamma^{\varepsilon}})\big{]},$
so that
(2.29) $\begin{split}&\mathbb{W}_{\alpha}(P_{t}^{*}\gamma^{\varepsilon+r},P_{t}^{*}\gamma^{\varepsilon})^{2}=\sup_{[f]_{\alpha}\leq 1}\big{|}P_{t}f(\gamma^{\varepsilon+r})-P_{t}f(\gamma^{\varepsilon})\big{|}^{2}\leq I_{1}+I_{2},\\ &I_{1}:=2\sup_{[f]_{\alpha}\leq 1}\bigg{|}\int_{\mathbb{R}^{d}}\mathbb{E}\left[f(X^{x,\gamma^{\varepsilon}}_{t})(R_{t}^{\varepsilon,r}-1)\right]\gamma^{\varepsilon+r}(\text{\rm{d}}x)\bigg{|}^{2},\\ &I_{2}:=2\bigg{|}\mathbb{E}\int_{0}^{r}\frac{\text{\rm{d}}}{\text{\rm{d}}\theta}P_{t}^{\gamma^{\varepsilon}}f(X_{0}^{\gamma^{\varepsilon+\theta}})\text{\rm{d}}\theta\bigg{|}^{2}=2\bigg{|}\mathbb{E}\int_{0}^{r}\Big{\{}\nabla_{X_{0}^{\tilde{\gamma}}-X_{0}^{\gamma}}P_{t}^{\gamma^{\varepsilon}}f(X_{0}^{\gamma^{\varepsilon+\theta}})\Big{\}}\text{\rm{d}}\theta\bigg{|}^{2}.\end{split}$
Below we estimate $I_{1}$ and $I_{2}$ respectively.
By (2.28), we obtain
(2.30) $\begin{split}&\mathbb{E}|R_{t}^{\varepsilon,r}-1|^{2}=\mathbb{E}\big{[}(R_{t}^{\varepsilon,r})^{2}-1\big{]}\leq\mathrm{esssup}_{\Omega}(\text{\rm{e}}^{\int_{0}^{t}|\eta_{s}^{\varepsilon,r}|^{2}\text{\rm{d}}s}-1)\\ &\leq\mathrm{esssup}_{\Omega}\left(\text{\rm{e}}^{\int_{0}^{t}|\eta_{s}^{\varepsilon,r}|^{2}\text{\rm{d}}s}\int_{0}^{t}|\eta_{s}^{\varepsilon,r}|^{2}\text{\rm{d}}s\right)\\ &\leq\psi(\varepsilon,r)\int_{0}^{t}\big{\{}\mathbb{W}_{\alpha}(P_{s}^{\ast}\gamma^{\varepsilon},P_{s}^{\ast}\gamma^{\varepsilon+r})^{2}+\mathbb{W}_{k}(P_{s}^{\ast}\gamma^{\varepsilon},P_{s}^{\ast}\gamma^{\varepsilon+r})^{2}\big{\}}\text{\rm{d}}s,\end{split}$
where for $c_{2}:=2c_{1}^{2}$,
(2.31) $\psi(\varepsilon,r):=c_{2}\text{\rm{e}}^{c_{2}\int_{0}^{T}\{\mathbb{W}_{\alpha}(P_{s}^{\ast}\gamma^{\varepsilon},P_{s}^{\ast}\gamma^{\varepsilon+r})^{2}+\mathbb{W}_{k}(P_{s}^{\ast}\gamma^{\varepsilon},P_{s}^{\ast}\gamma^{\varepsilon+r})^{2}\}\text{\rm{d}}s}.$
By (2.3) and (2.4) for $n=k$, we have
(2.32) $\bar{\psi}:=\sup_{\varepsilon,r\in[0,1]}\psi(\varepsilon,\gamma)<\infty.$
Combining this with (2.1), (2.15), (2.30) and (2.22) with $y=\tilde{\theta}^{\lambda,\gamma^{\varepsilon}}_{t}(x)$, where $\tilde{\theta}^{\lambda,\gamma^{\varepsilon}}_{t}(x)$ is defined in (2.12) with $\gamma^{\varepsilon}$ replacing $\gamma$, we can find constants $k_{1},k_{2}>1$ such that
$\bigg{(}\int_{\mathbb{R}^{d}}\sup_{[f]_{\alpha}\leq 1}\Big{|}\mathbb{E}\left[f(X^{x,\gamma^{\varepsilon}}_{t})(R_{t}^{\varepsilon,r}-1)\right]\Big{|}\gamma^{\varepsilon+r}(\text{\rm{d}}x)\bigg{)}^{2}$
$\leq\bigg{(}\int_{\mathbb{R}^{d}}\alpha\big{(}k_{1}t^{\frac{1}{2}}\big{)}\sup_{x}\big{(}\mathbb{E}[|R_{t}^{\varepsilon,r}-1|^{2}]\big{)}^{\frac{1}{2}}\gamma^{\varepsilon+r}(\text{\rm{d}}x)\bigg{)}^{2}$
$\leq\alpha\big{(}k_{1}t^{\frac{1}{2}}\big{)}^{2}\sup_{x}\mathbb{E}[|R_{t}^{\varepsilon,r}-1|^{2}]$
$\leq k_{2}\alpha\big{(}t^{\frac{1}{2}}\big{)}^{2}\psi(\varepsilon,r)\int_{0}^{t}\big{\{}\mathbb{W}_{\alpha}(P_{s}^{\ast}\gamma^{\varepsilon},P_{s}^{\ast}\gamma^{\varepsilon+r})^{2}+\mathbb{W}_{k}(P_{s}^{\ast}\gamma^{\varepsilon},P_{s}^{\ast}\gamma^{\varepsilon+r})^{2}\big{\}}\text{\rm{d}}s,\ \ t\in[0,T].$
Combining this with (2.1), (2.18), (2.26), (2.25), and letting
$\Gamma_{t}(\varepsilon,r):=\mathbb{W}_{\alpha}(P_{t}^{*}\gamma^{\varepsilon},P_{t}^{*}\gamma^{\varepsilon+r})^{2}+\int_{0}^{t}\mathbb{W}_{\alpha}(P_{s}^{*}\gamma^{\varepsilon},P_{s}^{*}\gamma^{\varepsilon+r})^{2}\text{\rm{d}}s,$
we find a constant $c_{4}>0$ such that
(2.33) $I_{1}\leq c_{4}\alpha\big{(}\sqrt{T}\big{)}^{2}\psi(\varepsilon,r)\bigg{(}r^{2}\mathbb{W}_{k}(\gamma,\tilde{\gamma})^{2}+\int_{0}^{t}\Gamma_{s}(\varepsilon,r)\text{\rm{d}}s\bigg{)},\ \ t\in[0,T].$
Combining this with Lemma 2.3(1), (2.15), (A) and (2.22) with $y=\tilde{\theta}^{\lambda,\gamma^{\varepsilon}}_{t}(x)$, we find constants $c_{5}>0$ such that
$\begin{split}\sup_{[f]_{\alpha}\leq 1}\big{|}\nabla P^{\gamma^{\varepsilon}}_{t}f\big{|}(x)&\leq\frac{c_{5}}{\sqrt{t}}\alpha\big{(}t^{\frac{1}{2}}\big{)}.\end{split}$
Combining this with (2.1), we find a constant $c_{6}>0$ such that
(2.34) $\begin{split}I_{2}&\leq 2\bigg{(}\mathbb{E}\bigg{[}|X_{0}^{\gamma}-X_{0}^{\tilde{\gamma}}|\int_{0}^{r}|\nabla P_{t}^{\gamma^{\varepsilon}}f(X_{0}^{\gamma^{\varepsilon+\theta}})|\text{\rm{d}}\theta\bigg{]}\bigg{)}^{2}\\ &\leq\frac{c_{6}\alpha\big{(}t^{\frac{1}{2}}\big{)}^{2}}{t}r^{2}(\mathbb{E}|X_{0}^{\gamma}-X_{0}^{\tilde{\gamma}}|)^{2}\\ &\leq\frac{c_{6}\alpha\big{(}t^{\frac{1}{2}}\big{)}^{2}r^{2}}{t}\big{(}\mathbb{E}[|X_{0}^{\gamma}-X_{0}^{\tilde{\gamma}}|^{k}]\big{)}^{\frac{2}{k}}.\end{split}$
Let
(2.35) $\tilde{\alpha}(r):=\bigg{(}\int_{0}^{r}\frac{\alpha(t)^{2}}{t}\text{\rm{d}}t\bigg{)}^{\frac{1}{2}},\ \ \ r\geq 0.$
By (2.1), we find some constant $c^{\prime}>0$ such that
(2.36) $\begin{split}&\int_{0}^{T}\frac{\alpha(rt^{\frac{1}{2}})^{2}}{t}\text{\rm{d}}t=2\int_{0}^{rT^{\frac{1}{2}}}\frac{\alpha(s)^{2}}{s}\text{\rm{d}}s\leq c^{\prime}\tilde{\alpha}(r)^{2}<\infty,\ \ r\geq 1.\end{split}$
So, (2.34) together with (2.29) and (2.33) yields that for some constant $c_{7}>0$,
(2.37) $\begin{split}&\Gamma_{t}(\varepsilon,r)\leq c_{7}r^{2}\mathbb{W}_{k}(\gamma,\tilde{\gamma})^{2}H_{t}(\varepsilon,r)+c_{7}\psi(\varepsilon,r)\int_{0}^{t}\Gamma_{s}(\varepsilon,r)\text{\rm{d}}s,\\ &H_{t}(\varepsilon,r):=\psi(\varepsilon,r)+\tilde{\alpha}\big{(}1\big{)}^{2}+\frac{\alpha\big{(}t^{\frac{1}{2}}\big{)}^{2}}{t},\ \ \ \varepsilon,r\in[0,1],t\in[0,T].\end{split}$
By Gronwall’s inequality and (2.37), for any $\varepsilon,r\in[0,1]$ we have
$\mathbb{W}_{\alpha}(P_{t}^{*}\gamma^{\varepsilon},P_{t}^{*}\gamma^{\varepsilon+r})^{2}\leq\Gamma_{t}(\varepsilon,r)$
$\leq c_{7}r^{2}\mathbb{W}_{k}(\gamma,\tilde{\gamma})^{2}\bigg{\{}H_{t}(\varepsilon,r)+c_{7}\psi(\varepsilon,r)\text{\rm{e}}^{c_{7}\psi(\varepsilon,r)T}\int_{0}^{t}H_{s}(\varepsilon,r)\text{\rm{d}}s\bigg{\}},\ \ t\in[0,T].$
This together with (2.18), (2.36) and (2.31)-(2.32) implies that $\psi(\varepsilon,r)$ is bounded in $(\varepsilon,r)\in[0,1]^{2}$ with $\psi(\varepsilon,r)\rightarrow c_{2}$ as $r\rightarrow 0$, so that by the dominated convergence theorem we find a constant $c>0$ such that
(2.38) $\begin{split}&\limsup_{r\downarrow 0}\frac{\mathbb{W}_{\alpha}(P_{t}^{*}\gamma^{\varepsilon},P_{t}^{*}\gamma^{\varepsilon+r})}{r}\leq c\mathbb{W}_{k}(\tilde{\gamma},\gamma)\bigg{\{}\frac{\alpha(t^{\frac{1}{2}})}{\sqrt{t}}+1\bigg{\}}.\end{split}$
By the triangle inequality,
$\big{|}\mathbb{W}_{\alpha}(P_{t}^{*}\gamma,P_{t}^{*}\gamma^{\varepsilon})-\mathbb{W}_{\alpha}(P_{t}^{*}\gamma,P_{t}^{*}\gamma^{\varepsilon+r})\big{|}\leq\mathbb{W}_{\alpha}(P_{t}^{*}\gamma^{\varepsilon},P_{t}^{*}\gamma^{\varepsilon+r}),\ \ \varepsilon,r\in[0,1],$
so that (2.38) implies that $\mathbb{W}_{\alpha}(P_{t}^{*}\gamma,P_{t}^{*}\gamma^{\varepsilon})$ is Lipschitz continuous (hence a.e. differentiable) in $\varepsilon\in[0,1]$ for any $t\in(0,T]$, and
$\Big{|}\frac{\text{\rm{d}}}{\text{\rm{d}}\varepsilon}\mathbb{W}_{\alpha}(P_{t}^{*}\gamma,P_{t}^{*}\gamma^{\varepsilon})\Big{|}\leq\limsup_{r\downarrow 0}\frac{\mathbb{W}_{\alpha}(P_{t}^{*}\gamma^{\varepsilon},P_{t}^{*}\gamma^{\varepsilon+r})}{r}\leq c\mathbb{W}_{k}(\tilde{\gamma},\gamma)\bigg{\{}\frac{\alpha(t^{\frac{1}{2}})}{\sqrt{t}}+1\bigg{\}},\ \ \varepsilon\in[0,1].$
Noting that $\gamma^{1}=\tilde{\gamma}$, this implies the desired estimate. ∎
Proof of Theorem 2.1.
Let $k=2$. According to [28, Theorem 2.5] for $D=\mathbb{R}^{d}$, see also [32, Theorem 4.1], (A) implies the following log-Harnack inequality for some constant $c_{0}>0$ and any $\gamma\in\mathscr{P}_{2}$:
$P_{t}^{\gamma}\log f(x)\leq\log P_{t}^{\gamma}f(y)+\frac{c_{0}}{t}|x-y|^{2},\ \ x,y\in\mathbb{R}^{d},t\in(0,T],f\in\mathscr{B}_{b}^{+}(\mathbb{R}^{d}).$
Then by [28, (4.13)], see also [13, Theorem 2.1], it suffices to find a constant $c>0$ such that
(2.39) $\begin{split}&\sup_{t\in(0,T]}\log\mathbb{E}[|R_{t}^{\gamma,\tilde{\gamma}}|^{2}]\leq c\mathbb{W}_{2}(\gamma,\tilde{\gamma})^{2},\ \ \gamma,\tilde{\gamma}\in\mathscr{P}_{2},\end{split}$
where
$R_{t}^{\gamma,\tilde{\gamma}}:=\text{\rm{e}}^{\int_{0}^{t}\langle\eta_{s}^{\gamma,\tilde{\gamma}},\text{\rm{d}}W_{s}\rangle-\frac{1}{2}\int_{0}^{t}|\eta_{s}^{\gamma,\tilde{\gamma}}|^{2}\text{\rm{d}}s},$
$\eta_{s}^{\gamma,\tilde{\gamma}}:=\big{\{}\sigma_{s}^{*}(\sigma_{s}\sigma_{s}^{*})^{-1}\big{\}}(X_{s}^{\gamma})\big{\{}b_{s}(X_{s}^{\gamma},P_{s}^{*}\tilde{\gamma})-b_{s}(X_{s}^{\gamma},P_{s}^{*}\gamma)\big{\}},\ \ s\leq t\leq T.$
Noting that (A) implies
$|\eta_{s}^{\gamma,\tilde{\gamma}}|^{2}\leq c_{1}\big{\{}\mathbb{W}_{\alpha}(P_{s}^{\ast}\gamma,P_{s}^{*}\tilde{\gamma})^{2}+\mathbb{W}_{2}(P_{s}^{\ast}\gamma,P_{s}^{*}\tilde{\gamma})^{2}\big{\}},\ \ s\in[0,T]$
for some constant $c_{1}>0$, we have
$\mathbb{E}[|R_{t}^{\gamma,\tilde{\gamma}}|^{2}]\leq\text{\rm{e}}^{c_{1}\int_{0}^{t}\{\mathbb{W}_{\alpha}(P_{s}^{\ast}\gamma,P_{s}^{*}\tilde{\gamma})^{2}+\mathbb{W}_{2}(P_{s}^{\ast}\gamma,P_{s}^{*}\tilde{\gamma})^{2}\}\text{\rm{d}}s}.$
Moreover, by (2.36) and Lemma 2.6, there exists a constant $c>0$ such that
$\sup_{t\in[0,T]}\int_{0}^{t}\big{\{}\mathbb{W}_{\alpha}(P_{s}^{\ast}\gamma,P_{s}^{*}\tilde{\gamma})^{2}+\mathbb{W}_{2}(P_{s}^{\ast}\gamma,P_{s}^{*}\tilde{\gamma})^{2}\big{\}}\text{\rm{d}}s\leq c\mathbb{W}_{2}(\gamma,\tilde{\gamma})^{2}.$
Therefore, (2.39) holds for some constant $c>0$.
3 Bismut Formula
Let $k\in(1,\infty)$ and denote $k^{*}:=\frac{k}{k-1}$. In this part, we establish the Bismut formula for the intrinsic derivative of $P_{t}f(\gamma)$ for $\gamma\in\mathscr{P}_{k}.$ To this end, we assume
(3.1) $b_{t}(x,\mu)=b_{t}^{(0)}(x)+B_{t}(x,\mu,\mu(V)),\ \ t\in[0,T],x\in\mathbb{R}^{d},\mu\in\mathscr{P},$
where for a Banach space $(\mathbb{B},\|\cdot\|_{\mathbb{B}})$,
$V:\mathbb{R}^{d}\rightarrow\mathbb{B},\ \ \ \ B:[0,T]\times\mathbb{R}^{d}\times\mathscr{P}\times\mathbb{B}\rightarrow\mathbb{R}^{d}$
are measurable such that $[V]_{\alpha}\leq 1$ for some $\alpha\in\mathscr{A}$, i.e. $V$ is only square root Dini continuous and hence $B_{t}(x,\mu,\mu(V))$ may be not intrinsically differentiable in $\mu$, which produces essential difficulty. The following definition is taken from [3] using the idea of [1].
Definition 3.1.
Let $f\in C(\mathscr{P}_{k};\mathbb{B})$ for a Banach space $\mathbb{B}$. The function $f$ is called intrinsically differentiable at a point $\mu\in\mathscr{P}_{k}$, if
$T_{\mu,k}:=L^{k}(\mathbb{R}^{d}\rightarrow\mathbb{R}^{d};\mu)\ni\phi\mapsto D_{\phi}^{I}f(\mu):=\lim_{\varepsilon\downarrow 0}\frac{f(\mu\circ(id+\varepsilon\phi)^{-1})-f(\mu)}{\varepsilon}\in\mathbb{B}$
is a well defined bounded linear operator. In this case, the norm of the intrinsic derivative $D^{I}f(\mu)$ is given by
$\|D^{I}f(\mu)\|_{L^{k^{*}}(\mu)}:=\sup_{\|\phi\|_{L^{k}(\mu)}\leq 1}\|D^{I}_{\phi}f(\mu)\|_{\mathbb{B}}.$
If moreover
$\lim_{\|\phi\|_{T_{\mu,k}}\downarrow 0}\frac{\|f(\mu\circ(id+\phi)^{-1})-f(\mu)-D_{\phi}^{I}f(\mu)\|_{\mathbb{B}}}{\|\phi\|_{T_{\mu,k}}}=0,$
then $f$ is called $L$-differentiable at $\mu$. The function $f$ is called intrinsically differentiable or $L$-differentiable on $\mathscr{P}_{k}$, if it is so at any $\mu\in\mathscr{P}_{k}$.
Recall that a real function $f$ on a Banach space $\mathbb{B}$ is called Gateaux differentiable, if for any $z\in\mathbb{B}$,
$v\mapsto\nabla_{v}^{\mathbb{B}}f(z):=\lim_{\varepsilon\downarrow 0}\frac{f(z+\varepsilon v)-f(z)}{\varepsilon}$
is a well-defined bounded linear functional. In this case we denote
$\|\nabla^{\mathbb{B}}f(z)\|_{\mathbb{B}^{*}}:=\sup_{\|v\|_{\mathbb{B}}\leq 1}|\nabla_{v}^{\mathbb{B}}f(z)|.$
Moreover, $f$ is called Fréchet differentiable if it is Gateaux differentiable and
$\lim_{\|v\|_{\mathbb{B}}\downarrow 0}\frac{|f(z+v)-f(z)-\nabla_{v}^{\mathbb{B}}f(z)|}{\|v\|_{\mathbb{B}}}=0,\ \ z\in\mathbb{B}.$
It is well-known that a Gateaux differentiable function $f$ is Fréchet differentiable provided $\nabla_{v}^{\mathbb{B}}f(z)$ is continuous in $(v,z)\in\mathbb{B}\times\mathbb{B}$.
When $\mathbb{B}=\mathbb{R}^{l}$ for some $l\geq 1$, we simply denote $\nabla^{\mathbb{B}}=\nabla$.
3.1 Main result
We will establish a Bismut formula for the intrinsic derivative of $P_{t}f$ under the following assumption.
  1. (B)
    Let $k\in(1,\infty)$ and let $b$ in (3.1).
  2. $(B_{1})$
    $b^{(0)}$ and $\sigma$ satisfy the corresponding conditions in (A).
  3. $(B_{2})$
    There exists $\alpha\in\mathscr{A}$ and $K_{V}\geq 0$ such that
    $[V]_{\alpha}:=\sup_{x\neq y}\frac{\|V(x)-V(y)\|_{\mathbb{B}}}{\alpha(|x-y|)}\leq 1,\ \ \|V(x)\|_{\mathbb{B}}\leq K_{V}(1+|x|^{k-1}),\ \ x\in\mathbb{R}^{d}.$
  4. $(B_{3})$
    For any $t\in[0,T]$, we have $B_{t}\in C(\mathbb{R}^{d}\times\mathscr{P}_{k}\times\mathbb{B})$. $B_{t}(x,\mu,z)$ is differentiable in $x$, $L$-differentiable in $\mu\in\mathscr{P}_{k}$, and Fréchet differentiable in $z\in\mathbb{B}$ such that $\nabla_{v}^{\mathbb{B}}B_{t}(x,\mu,z)$ is continuous in $(v,z)\in\mathbb{B}\times\mathbb{B}$. Next, there exists a constant $K>0$ such that
    $|B_{t}(0,\delta_{0},0)|\leq K,$
    $|\nabla B_{t}(\cdot,\mu,z)(x)|+\|D^{L}B_{t}(x,\cdot,z)(\mu)\|_{L^{k^{*}}(\mu)}+\|\nabla^{\mathbb{B}}B_{t}(x,\mu,\cdot)(z)\|_{\mathbb{B}^{*}}\leq K,$
    $\ \ (t,x,\mu,z)\in[0,T]\times\mathbb{R}^{d}\times\mathscr{P}_{k}\times\mathbb{B}.$
    Moreover, for any $(t,x,\mu,z)\in[0,T]\times\mathbb{R}^{d}\times\mathscr{P}_{k}\times\mathbb{B}$, there exists a constant $c(t,x,\mu,z)>0$ such that
    |$|\{D^{L}B_{t}(x,\cdot,z)(\mu)(y)|\leq c(t,x,\mu,z)(1+|y|^{k-1}),\ \ y\in\mathbb{R}^{d}.$
Example 3.1.
Let $l$ be a positive integer and let $\alpha(s)=s^{\varepsilon}$ for some $\varepsilon\in(0,1\wedge(k-1))$. Then $V:\mathbb{R}^{d}\rightarrow\mathbb{R}^{l}$ with
$|V(x)-V(y)|\leq\alpha(|x-y|),\ \ x,y\in\mathbb{R}^{d}$
satisfies ($B_{2}$). Moreover, let $g\in C^{1}(\mathbb{R}^{d}\times\mathbb{R}^{d}\times\mathbb{R}^{l};\mathbb{R}^{d})$ with bounded gradient. Then $B_{t}(x,\mu,z)=\int_{\mathbb{R}^{d}}g(x,y,z)\mu(\text{\rm{d}}y)$ satisfies ($B_{3}$).
Since (B) implies (A), as explained before that under this assumption (1.1) is well-posed for distributions in $\mathscr{P}_{k}.$
For $\mu\in\mathscr{P}_{k}$, consider the decoupled SDE
(3.2) $\begin{split}&\text{\rm{d}}X_{t}^{x,\mu}=\big{\{}b_{t}^{(0)}(X_{t}^{x,\mu})+B_{t}(X_{t}^{x,\mu},P_{t}^{*}\mu,P_{t}V(\mu))\big{\}}\text{\rm{d}}t+\sigma_{t}(X_{t}^{x,\mu})\text{\rm{d}}W_{t},\\ &\qquad\qquad\qquad X_{0}^{x,\mu}=x,\ \ t\in[0,T].\end{split}$
To state the Bismut formula for $P_{t}f$, we introduce the quantity $I_{t}^{f}$: for fixed $t\in(0,T]$, let
(3.3) $\begin{split}I_{t}^{f}(\mu,\phi)&:=\frac{1}{t}\int_{\mathbb{R}^{d}}\mathbb{E}\bigg{[}f(X_{t}^{x,\mu})\int_{0}^{t}\big{\langle}\zeta_{s}(X_{s}^{x,\mu})\nabla_{\phi(x)}X_{s}^{x,\mu},\text{\rm{d}}W_{s}\big{\rangle}\bigg{]}\mu(\text{\rm{d}}x),\\ \zeta_{s}&:=\sigma_{s}^{*}(\sigma_{s}\sigma_{s}^{*})^{-1},\ \ \ s\in[0,t],\mu\in\mathscr{P}_{k},\phi\in T_{\mu,k}.\end{split}$
By $(B_{1})$ and (2.14), we find a constant $c>0$ such that
(3.4) $|I_{t}^{f}(\mu,\phi)|\leq\frac{c}{\sqrt{t}}\big{(}P_{t}|f|^{k^{*}}(\mu)\big{)}^{\frac{1}{k^{*}}}\|\phi\|_{L^{k}(\mu)},\ \ \mu\in\mathscr{P}_{k},\phi\in T_{\mu,k}.$
Next, let $X_{0}^{\mu}$ be $\mathscr{F}_{0}$-measurable such that $\mathscr{L}_{X_{0}^{\mu}}=\mu$, and let $X_{t}^{\mu}$ solve (1.1) with initial value $X_{0}^{\mu}$. For any $\varepsilon\geq 0$, denote
$\mu_{\varepsilon}:=\mu\circ(id+\varepsilon\phi)^{-1},\ \ X_{0}^{\mu_{\varepsilon}}:=X_{0}^{\mu}+\varepsilon\phi(X_{0}^{\mu}).$
Let $X_{t}^{\mu_{\varepsilon}}$ solve (1.1) with initial value $X_{0}^{\mu_{\varepsilon}}$. So,
$X_{t}^{\mu}=X_{t}^{\mu_{0}},\ \ P_{t}^{*}\mu_{\varepsilon}=\mathscr{L}_{X_{t}^{\mu_{\varepsilon}}},\ \ t\in[0,T],\varepsilon\geq 0.$
Theorem 3.2.
[…]
  1. $(1)$
[…]
  2. $(2)$
    The limit
    (3.5) $\nabla_{\phi}X_{t}^{\mu}:=\lim_{\varepsilon\downarrow 0}\frac{X_{t}^{\mu_{\varepsilon}}-X_{t}^{\mu}}{\varepsilon},\ \ \ \ t\in[0,T]$
    exists in $L^{k}(\Omega\rightarrow C([0,T],\mathbb{R}^{d}),\mathbb{P})$, and there exists a constant $c>0$ such that
    $\mathbb{E}\Big{[}\sup_{t\in[0,T]}|\nabla_{\phi}X_{t}^{\mu}|^{k}\Big{]}\leq c\|\phi\|_{L^{k}(\mu)}^{k},\ \ \mu\in\mathscr{P}_{k}.$
  3. $(3)$
    For any $t\in(0,T]$ and $f\in\mathscr{B}_{b}(\mathbb{R}^{d})$, $P_{t}f$ is intrinsically differentiable on $\mathscr{P}_{k}$. Moreover, for any $\mu\in\mathscr{P}_{k}$ and $\phi\in T_{\mu,k}$,
    (3.6) $\begin{split}&D_{\phi}^{I}P_{t}f(\mu)=I_{t}^{f}(\mu,\phi)+\mathbb{E}\bigg{[}f(X_{t}^{\mu})\int_{0}^{t}\Big{\langle}\zeta_{s}(X_{s}^{\mu})\big{\{}N_{s}+\tilde{N}_{s}\big{\}},\ \text{\rm{d}}W_{s}\Big{\rangle}\bigg{]},\\ &N_{s}:=\big{\{}\nabla^{\mathbb{B}}_{D_{\phi}^{I}P_{s}V(\mu)}B_{s}(X_{s}^{\mu},P_{s}^{\ast}\mu,\cdot)\big{\}}(P_{s}V(\mu)),\\ &\tilde{N}_{s}:=\mathbb{E}\big{[}\big{\langle}\{D^{L}B_{s}(y,\cdot,P_{s}V(\mu))\}(P_{s}^{\ast}\mu)(X_{s}^{\mu}),\nabla_{\phi}X_{s}^{\mu}\big{\rangle}\big{]}|_{y=X_{s}^{\mu}},\end{split}$
[…]
By (3.4)-(3.6) and ($B_{3}$), we find a constant $c>0$ such that
$\|D^{I}P_{t}f(\mu)\|_{L^{k^{*}}(\mu)}\leq c\frac{\{P_{t}|f|^{k^{*}}(\mu)\}^{\frac{1}{k^{*}}}}{\sqrt{t}},\ \ t\in(0,T],f\in\mathscr{B}_{b}(\mathbb{R}^{d}),\mu\in\mathscr{P}_{k}.$
Remark 3.3.
Note that if $\nabla V$ is continuous and bounded, then the chain rule (see for instance [19, Proposition 3.1]) implies
(3.7) $D^{I}_{\phi}P_{t}V(\mu)=\mathbb{E}\nabla_{\nabla_{\phi}X_{t}^{\mu}}V(X_{t}^{\mu}).$
However, in the present case, $V$ only satisfies $(B_{2})$ so that (3.7) does not hold and there is essential difficulty in the study of $D^{I}_{\phi}P_{t}V$.
To explain the main idea in the proof of Theorem 3.2, we first figure out a sketch. By the definition of the intrinsic derivative, we intend to calculate
(3.8) $D_{\phi}^{I}P_{t}f(\mu):=\lim_{\varepsilon\downarrow 0}\frac{P_{t}f(\mu_{\varepsilon})-P_{t}f(\mu)}{\varepsilon}=\lim_{\varepsilon\downarrow 0}\frac{\mathbb{E}[f(X_{t}^{\mu_{\varepsilon}})-f(X_{t}^{\mu})]}{\varepsilon}.$
To this end, for any $\gamma\in\mathscr{P}_{k},$ consider the decoupled SDE (3.2) for $\gamma$ replacing $\mu$, i.e.
(3.9) $\begin{split}\text{\rm{d}}X_{t}^{x,\gamma}=\big{\{}b_{t}^{(0)}(X_{t}^{x,\gamma})+&B_{t}\big{(}X_{t}^{x,\gamma},P_{t}^{\ast}\gamma,P_{t}V(\gamma)\big{)}\big{\}}\text{\rm{d}}t+\sigma_{t}(X_{t}^{x,\gamma})\text{\rm{d}}W_{t},\\ &t\in[0,T],X_{0}^{x,\gamma}=x\in\mathbb{R}^{d}.\end{split}$
Let
$P_{t}^{\gamma}f(x):=\mathbb{E}[f(X_{t}^{x,\gamma})],\ \ \ x\in\mathbb{R}^{d},$
$P_{t}^{\gamma}f(\tilde{\gamma}):=\int_{\mathbb{R}^{d}}P_{t}^{\gamma}f\text{\rm{d}}\tilde{\gamma},\ \ t\geq 0,f\in\mathscr{B}_{b}(\mathbb{R}^{d}),\gamma,\tilde{\gamma}\in\mathscr{P}_{k}.$
For $\varepsilon\geq 0$, let $X_{t}^{\mu_{\varepsilon},\gamma}$ be the solution of (3.9) with initial value $X_{0}^{\varepsilon}$, i.e,
$\begin{split}\text{\rm{d}}X_{t}^{\mu_{\varepsilon},\gamma}=\big{\{}b_{t}^{(0)}(X_{t}^{\mu_{\varepsilon},\gamma})+&B_{t}\big{(}X_{t}^{\mu_{\varepsilon},\gamma},P_{t}^{\ast}\gamma,P_{t}V(\gamma)\big{)}\big{\}}\text{\rm{d}}t+\sigma_{t}(X_{t}^{\mu_{\varepsilon},\gamma})\text{\rm{d}}W_{t},\\ &t\in[0,T],\ X_{0}^{\mu_{\varepsilon},\gamma}=X_{0}^{\varepsilon}.\end{split}$
Then $X_{t}^{\mu_{\varepsilon},\mu_{\varepsilon}}$ solves (1.1) with initial value $X_{0}^{\varepsilon}$, so that
$P_{t}f(\mu_{\varepsilon})=P_{t}^{\mu_{\varepsilon}}f(\mu_{\varepsilon})=\mathbb{E}[f(X_{t}^{\mu_{\varepsilon},\mu_{\varepsilon}})],\ \ \varepsilon\geq 0,t\in[0,T],f\in\mathscr{B}_{b}(\mathbb{R}^{d}).$
Noting that $\mu_{0}=\mu$, (3.8) reduces to
(3.10) $\begin{split}&D_{\phi}^{I}P_{t}f(\mu)=\lim_{\varepsilon\downarrow 0}\frac{P_{t}^{\mu_{\varepsilon}}f(\mu_{\varepsilon})-P_{t}^{\mu}f(\mu)}{\varepsilon}\\ &=\lim_{\varepsilon\downarrow 0}\Big{\{}\frac{P_{t}^{\mu}f(\mu_{\varepsilon})-P_{t}^{\mu}f(\mu)}{\varepsilon}+\frac{P_{t}^{\mu_{\varepsilon}}f(\mu_{\varepsilon})-P_{t}^{\mu}f(\mu_{\varepsilon})}{\varepsilon}\Big{\}}.\end{split}$
So, to calculate $D_{\phi}^{I}P_{t}f(\mu),$ we only need to study the limits of
$J_{1}f(t,\varepsilon):=\frac{P_{t}^{\mu}f(\mu_{\varepsilon})-P_{t}^{\mu}f(\mu)}{\varepsilon},\ \ J_{2}f(t,\varepsilon):=\frac{P_{t}^{\mu_{\varepsilon}}f(\mu_{\varepsilon})-P_{t}^{\mu}f(\mu_{\varepsilon})}{\varepsilon}.$
By Lemma 2.3(1), for any $t\in(0,T],\varepsilon\geq 0$ and $f\in\mathscr{B}_{b}(\mathbb{R}^{d})$, we have
(3.11) $\begin{split}&\frac{\text{\rm{d}}}{\text{\rm{d}}\varepsilon}P_{t}^{\mu}f(\mu_{\varepsilon}):=\lim_{r\downarrow 0}\frac{P_{t}^{\mu}f(\mu_{\varepsilon+r})-P_{t}^{\mu}f(\mu_{\varepsilon})}{r}\\ &=\int_{\mathbb{R}^{d}}\mathbb{E}\bigg{[}f(X_{t}^{x+\varepsilon\phi(x),\mu})\frac{1}{t}\int_{0}^{t}\big{\langle}\zeta_{s}(X_{s}^{x+\varepsilon\phi(x),\mu})\nabla_{\phi(x)}X_{s}^{x+\varepsilon\phi(x),\mu},\text{\rm{d}}W_{s}\big{\rangle}\bigg{]}\mu(\text{\rm{d}}x)\\ &=:I_{t}^{f}(\varepsilon,\mu,\phi),\ \ t\in(0,T].\end{split}$
In particular,
(3.12) $\lim_{\varepsilon\downarrow 0}J_{1}f(t,\varepsilon)=\lim_{\varepsilon\downarrow 0}\frac{P_{t}^{\mu}f(\mu_{\varepsilon})-P_{t}^{\mu}f(\mu)}{\varepsilon}=I_{t}^{f}(\mu,\phi),\ \ t\in(0,T].$
Consequently, it remains to prove
$\lim_{\varepsilon\rightarrow 0}J_{2}f(t,\varepsilon)=\mathbb{E}\bigg{[}f(X_{t}^{\mu})\int_{0}^{t}\Big{\langle}\zeta_{s}(X_{s}^{\mu})\big{\{}N_{s}+\tilde{N}_{s}\big{\}},\ \text{\rm{d}}W_{s}\Big{\rangle}\bigg{]}$
for $N_{s},\tilde{N}_{s}$ defined in (3.6), which involves in $D^{I}P_{t}V(\mu)$ and $\nabla_{\phi}X_{t}^{\mu}$ defined in (3.5). Therefore, we will first study $D^{I}P_{t}V(\mu)$ and $\nabla_{\phi}X_{t}^{\mu}$.
To this end, let $u_{t}=u_{t}^{\lambda_{0},\mu},\Theta_{t}=\Theta_{t}^{\lambda_{0},\mu}$ be defined in (2.7) and (2.10) for large enough $\lambda_{0}>0$ such that (2.9) holds for $\lambda=\lambda_{0}$, and let
$\mathscr{C}_{k}:=\left\{\tilde{v}\in L^{k}\big{(}\Omega\rightarrow C([0,T];\mathbb{R}^{d}),\mathbb{P}\big{)}\ \text{is\ }\mathscr{F}_{t}\text{-adapted,\ }\tilde{v}_{0}=\nabla_{\phi(X_{0}^{\mu})}\Theta_{0}(X_{0}^{\mu})\right\},$
which is complete under the metric
$\rho_{\delta}(\tilde{v}^{1},\tilde{v}^{2}):=\left(\mathbb{E}\sup_{t\in[0,T]}\text{\rm{e}}^{-\delta kt}|\tilde{v}_{t}^{1}-\tilde{v}^{2}_{t}|^{k}\right)^{\frac{1}{k}},\ \ \tilde{v}^{1},\tilde{v}^{2}\in\mathscr{C}_{k},\delta>0.$
For any $s\in[0,T]$, $\eta\in L^{k}(\Omega\rightarrow\mathbb{R}^{d},\mathbb{P})$, define
$\psi_{s}(v):=\mathbb{E}\big{[}\big{\langle}\{D^{L}B_{s}(y,\cdot,P_{s}V(\mu))\}(P_{s}^{*}\mu)(X_{s}^{\mu}),(\nabla\Theta_{s}(X_{s}^{\mu}))^{-1}\eta\big{\rangle}\big{]}|_{y=X_{s}^{\mu}}.$
Recall that $\tilde{\alpha}$ is defined in (2.35). For any $(v,\tilde{v})\in C([0,T];\mathbb{B})\times\mathscr{C}_{k}$, we denote
(3.13) $\begin{split}&N_{s}(v):=\nabla_{v_{s}}^{\mathbb{B}}B_{s}(X_{s}^{\mu},P_{s}^{*}\mu,\cdot)(P_{s}V(\mu)),\ \ \tilde{N}_{s}(\tilde{v}):=\psi_{s}(\tilde{v}_{s}),\end{split}$
and write
$\begin{split}&\tilde{I}_{t}^{V}(\mu,v,\tilde{v})=\tilde{I}_{t}^{V,1}(\mu,v,\tilde{v})+\tilde{I}_{t}^{V,2}(\mu,v,\tilde{v}),\\ &\tilde{I}_{t}^{V,1}(\mu,v,\tilde{v}):=\mathbb{E}\bigg{[}V(X_{t}^{\mu})\int_{0}^{t}\frac{\alpha(s^{\frac{1}{2}})}{\{\tilde{\alpha}(s^{\frac{1}{2}})s\}^{\frac{1}{2}}}\Big{\langle}\zeta_{s}(X_{s}^{\mu})N_{s}(v),\ \text{\rm{d}}W_{s}\Big{\rangle}\bigg{]},\\ &\tilde{I}_{t}^{V,2}(\mu,v,\tilde{v}):=\mathbb{E}\bigg{[}V(X_{t}^{\mu})\int_{0}^{t}\Big{\langle}\zeta_{s}(X_{s}^{\mu})\tilde{N}_{s}(\tilde{v}),\ \text{\rm{d}}W_{s}\Big{\rangle}\bigg{]},\ \ t\in[0,T].\end{split}$
Consider the following equation for $(v,\tilde{v})\in C([0,T];\mathbb{B})\times\mathscr{C}_{k}$:
(3.14) $\begin{split}&\text{\rm{d}}\tilde{v}_{t}=\Big{[}\nabla_{(\nabla\Theta_{t}(X_{t}^{\mu}))^{-1}\tilde{v}_{t}}\big{\{}B_{t}(\cdot,P_{t}^{*}\mu,P_{t}V(\mu))+\lambda_{0}u_{t}\big{\}}(X_{t}^{\mu})\Big{]}\text{\rm{d}}t\\ &+\nabla_{\psi_{t}(\tilde{v}_{t})}\Theta_{t}(X_{t}^{\mu})\text{\rm{d}}t+{\frac{\alpha(t^{\frac{1}{2}})}{\{\tilde{\alpha}(t^{\frac{1}{2}})t\}^{\frac{1}{2}}}}\nabla_{N_{t}(v)}\Theta_{t}(X_{t}^{\mu})\text{\rm{d}}t\\ &+\Big{[}\nabla_{(\nabla\Theta_{t}(X_{t}^{\mu}))^{-1}\tilde{v}_{t}}\big{\{}(\nabla\Theta_{t})\sigma_{t}\big{\}}(X^{\mu}_{t})\Big{]}\,\text{\rm{d}}W_{t},\ \ \tilde{v}_{0}=\nabla_{\phi(X_{0}^{\mu})}\Theta_{0}(X_{0}^{\mu}),\\ &v_{t}={\frac{\{t\tilde{\alpha}(t^{\frac{1}{2}})\}^{\frac{1}{2}}}{\alpha(t^{\frac{1}{2}})}}\big{\{}I_{t}^{V}(\mu,\phi)+\tilde{I}_{t}^{V}(\mu,v,\tilde{v})\big{\}},\ \ t\in[0,T].\end{split}$
In the following two subsections, we prove the well-posedness of (3.14) and the formulas
(3.15) $\begin{split}&D_{\phi}^{I}P_{t}V(\mu)={\frac{\alpha(t^{\frac{1}{2}})}{\{t\tilde{\alpha}(t^{\frac{1}{2}})\}^{\frac{1}{2}}}}\,v_{t},\ \ t\in(0,T],\\ &\nabla_{\phi}X_{t}^{\mu}:=\lim_{\varepsilon\downarrow 0}\frac{X_{t}^{\mu_{\varepsilon}}-X_{t}^{\mu}}{\varepsilon}=(\nabla\Theta_{t}(X_{t}^{\mu}))^{-1}\tilde{v}_{t},\ \ \ t\in[0,T].\end{split}$
3.2 Well-posedness of (3.14)
Lemma 3.4.
[…]
(3.16) $\sup_{\|\phi\|_{L^{k}(\mu)}\leq 1}\|v_{t}(\mu,\phi)\|_{\mathbb{B}}\leq{c\,\sqrt{\tilde{\alpha}(t^{\frac{1}{2}})}},\ \ \mu\in\mathscr{P}_{k},t\in[0,T],$
and for any $j\geq 1$, there exists a constant $\tilde{c}>0$ such that
(3.17) $\mathbb{E}\Big{[}\sup_{t\in[0,T]}|\tilde{v}_{t}(\mu,\phi)|^{j}|\mathscr{F}_{0}\Big{]}\leq\tilde{c}|\phi(X_{0}^{\mu})|^{j}+{\tilde{c}\|\phi\|_{L^{k}(\mu)}^{j}},\ \ \mu\in\mathscr{P}_{k},t\in[0,T].$
Proof.
By Jensen’s inequality, for estimate (3.17) we only need to consider $j\geq k.$
As in the proof of [29, Lemma 4.1], there exists a constant $\theta>1$ and $0<f\in\mathscr{B}([0,T]\times\mathbb{R}^{d})$ such that
(3.18) [$\big{\|}\nabla\big{[}\{\nabla\Theta\}\sigma\big{]}\big{\|}\leq f,\ \ \ \mathbb{E}\int_{0}^{T}f_{s}(X_{s}^{\mu})^{2\theta}\text{\rm{d}}s<\infty.$
For any $n\geq 0$, define
$\tau_{n}:=T\wedge\inf\left\{t\geq 0:\int_{0}^{t}f_{s}(X_{s}^{\mu})^{2\theta}\text{\rm{d}}s\geq n\right\},$
$\mathscr{C}_{k,n}:=\big{\{}(z_{\cdot\wedge\tau_{n}}),\ \ z\in\mathscr{C}_{k}\big{\}}.$
We have $\lim_{n\rightarrow\infty}\tau_{n}=T$. Let
$\mathscr{V}_{0}:=\big{\{}v\in C([0,T];\mathbb{B}):\ v_{0}=0\big{\}}.$
Recall that $\tilde{\alpha}$ is defined in (2.35). By (3.18), $\|\nabla B\|_{\infty}<\infty$ due to (B), and
(3.19) $\int_{0}^{t}\frac{\alpha(rs^{\frac{1}{2}})^{2}}{s\tilde{\alpha}(rs^{\frac{1}{2}})}\text{\rm{d}}s=2\int_{0}^{rt^{\frac{1}{2}}}\frac{\alpha(s)^{2}}{s\tilde{\alpha}(s)}\text{\rm{d}}s=4\int_{0}^{rt^{\frac{1}{2}}}\tilde{\alpha}^{\prime}(s)\text{\rm{d}}s=4\tilde{\alpha}(rt^{\frac{1}{2}}),\ \ r\geq 0,$
we see that for any $(v,\tilde{v})\in\mathscr{V}_{0}\times\mathscr{C}_{k},$
$H_{s}(v,\tilde{v})$ $:=\tilde{v}_{0}+\int_{0}^{s}\nabla_{\psi_{t}(\tilde{v}_{t})}\Theta_{t}(X_{t}^{\mu})\text{\rm{d}}t$
$+\int_{0}^{s}\Big{\{}\nabla_{(\nabla\Theta_{t}(X_{t}^{\mu}))^{-1}\tilde{v}_{t}}\{B_{t}(\cdot,P_{t}^{*}\mu,P_{t}V(\mu))+\lambda_{0}u_{t}\}\Big{\}}(X_{t}^{\mu})\text{\rm{d}}t$
$+\int_{0}^{s}{\frac{\alpha(t^{\frac{1}{2}})}{\{\tilde{\alpha}(t^{\frac{1}{2}})t\}^{\frac{1}{2}}}}\nabla_{N_{t}(v)}\Theta_{t}(X_{t}^{\mu})\text{\rm{d}}t$
$+\int_{0}^{s}\nabla_{(\nabla\Theta_{t}(X_{t}^{\mu}))^{-1}\tilde{v}_{t}}\big{\{}(\nabla\Theta_{t})\sigma_{t}\big{\}}(X^{\mu}_{t})\,\text{\rm{d}}W_{t},\ \ s\in[0,T]$
is a continuous adapted process on $\mathbb{R}^{d}$ such that
$H_{\cdot\land\tau_{n}}:\mathscr{V}_{0}\times\mathscr{C}_{k,n}\rightarrow\mathscr{C}_{k,n},\ \ n\geq 1.$
On the other hand, let
(3.20) $\begin{split}&\bar{H}_{t}(v,\tilde{v}):={\frac{\{t\tilde{\alpha}(t^{\frac{1}{2}})\}^{\frac{1}{2}}}{\alpha(t^{\frac{1}{2}})}}\big{\{}I_{t}^{V}(\mu,\phi)+\tilde{I}_{t}^{V}(\mu,v,\tilde{v})\big{\}}.\end{split}$
We claim that it suffices to prove
  1. (i)
    for each $n\geq 1$, the map
    $\Phi^{(n)}:=(\bar{H},H_{\cdot\land\tau_{n}}):\ \mathscr{V}_{0}\times\mathscr{C}_{k,n}\rightarrow\mathscr{V}_{0}\times\mathscr{C}_{k,n}$
    is well-defined and has a unique fixed point $(v^{(n)},{\tilde{v}}^{(n)})$.
  2. (ii)
    there exists a constant $c>0$ such that
    $\sup_{\|\phi\|_{L^{k}(\mu)}\leq 1}\|v_{t}^{(n)}\|_{\mathbb{B}}\leq c{\,\sqrt{\tilde{\alpha}(t^{\frac{1}{2}})}},\ \ (t,\mu)\in[0,T]\times\mathscr{P}_{k},n\geq 1,$
    and for any $j\geq k$, there exists a constant $\tilde{c}>0$ such that
    $\mathbb{E}\Big{[}\sup_{t\in[0,T]}|{\tilde{v}}_{t}^{(n)}|^{j}|\mathscr{F}_{0}\Big{]}\leq\tilde{c}|\phi(X_{0}^{\mu})|^{j}+\tilde{c}{\|\phi\|_{L^{k}(\mu)}^{j}},\ \ (t,\mu)\in[0,T]\times\mathscr{P}_{k},n\geq 1.$
Indeed, by (i), the fixed points are consistent:
$\big{(}v^{(n)},{\tilde{v}}^{(n)}\big{)}_{t\land\tau_{n}}=\big{(}v^{(n+i)},{\tilde{v}}^{(n+i)}\big{)}_{t\land\tau_{n}},\ \ n,i\geq 1,t\in[0,T],$
so that $\mathbb{P}$-a.s.
$\big{(}v(\mu,\phi),\tilde{v}(\mu,\phi)\big{)}:=\lim_{n\rightarrow\infty}\big{(}v^{(n)},{\tilde{v}}^{(n)}\big{)}_{\cdot\land\tau_{n}}=\sum_{n=1}^{\infty}1_{[\tau_{n-1},\tau_{n})}\big{(}v^{(n)},{\tilde{v}}^{(n)}\big{)}_{\cdot\land\tau_{n}}$
exists in $C([0,T];\mathbb{B}\times\mathbb{R}^{d})$, and by (ii) the estimates (3.16) and (3.17) hold. Combining this with
(3.21) $\Phi^{(n)}\big{(}v^{(n)},{\tilde{v}}^{(n)}\big{)}:=\big{(}\bar{H},H_{\cdot\land\tau_{n}}\big{)}\big{(}v^{(n)},{\tilde{v}}^{(n)}\big{)}=\big{(}v^{(n)},{\tilde{v}}^{(n)}\big{)},\ \ n\geq 1,$
we see that $(v(\mu,\phi),\tilde{v}(\mu,\phi))$ solves (3.14), and the uniqueness follows from that of the fixed point for $\Phi^{(n)},n\geq 1.$ Next, we will prove (i) and (ii) one by one.
(1) Proof of (i).
(a) We first verify
(3.22) $\bar{H}:\mathscr{V}_{0}\times\mathscr{C}_{k}\rightarrow\mathscr{V}_{0},$
so that $\Phi^{(n)}:\mathscr{V}_{0}\times\mathscr{C}_{k,n}\rightarrow\mathscr{V}_{0}\times\mathscr{C}_{k,n}$ for every $n\geq 1$.
By (3.3), we have
(3.23) $I_{t}^{V}(\mu,\phi)=\frac{1}{t}\int_{\mathbb{R}^{d}}\mathbb{E}\bigg{[}\big{\{}V(X_{t}^{x,\mu})\big{\}}\int_{0}^{t}\big{\langle}\zeta_{s}(X_{s}^{x,\mu})\nabla_{\phi(x)}X_{s}^{x,\mu},\text{\rm{d}}W_{s}\big{\rangle}\bigg{]}\mu(\text{\rm{d}}x).$
Similar to (2.11)-(2.13), for any $\lambda\geq\lambda_{0}$, $\gamma\in\mathscr{P}_{k}$ and $x\in\mathbb{R}^{d}$, consider
$\text{\rm{d}}\theta^{\lambda,\gamma,B}_{t}(x)=B_{t}((\Theta^{\lambda,\gamma}_{t})^{-1}(\theta^{\lambda,\gamma,B}_{t}(x)),P_{t}^{\ast}\gamma,P_{t}V(\gamma))\text{\rm{d}}t,\ \ t\in[0,T],\theta^{\lambda,\gamma,B}_{0}(x)=\Theta^{\lambda,\gamma}_{0}(x),$
and let
$\tilde{\theta}^{\lambda,\gamma,B}_{t}(x)=(\Theta^{\lambda,\gamma}_{t})^{-1}(\theta^{\lambda,\gamma,B}_{t}(x)),\ \ t\in[0,T],x\in\mathbb{R}^{d}.$
Then we have
$\text{\rm{d}}\Theta^{\lambda,\gamma}_{t}(\tilde{\theta}^{\lambda,\gamma,B}_{t}(x))=B_{t}(\tilde{\theta}^{\lambda,\gamma,B}_{t}(x),P_{t}^{\ast}\gamma,P_{t}V(\gamma))\text{\rm{d}}t,\ \ t\in[0,T],\tilde{\theta}^{\lambda,\gamma,B}_{0}(x)=x\in\mathbb{R}^{d}.$
Repeating the proof of Lemma 2.3, we conclude that for any $p\geq 1$, there exists a constant $c_{p}>0$ such that
(3.24) $\mathbb{E}[|X^{x,\gamma}_{t}-\tilde{\theta}^{\lambda,\gamma,B}_{t}(x)|^{p}]\leq c_{p}t^{\frac{p}{2}},\ \ t\in[0,T],x\in\mathbb{R}^{d},\gamma\in\mathscr{P}_{k},$
and
(3.25) $\mathbb{E}\Big{[}\sup_{t\in[0,T]}|\nabla_{v}X_{t}^{x,\gamma}|^{p}\Big{]}\leq c_{p}|v|^{p},\ \ v\in\mathbb{R}^{d},\gamma\in\mathscr{P}_{k},x\in\mathbb{R}^{d}.$
By $[V]_{\alpha}\leq 1$, (2.22) in Lemma 2.5 for $p=2$ and $y=\tilde{\theta}^{\lambda_{0},\mu,B}_{t}(x)$, (3.25), (2.1), (3.24) and (3.23), we find a constant $c_{1}>0$ such that
(3.26) $\begin{split}\|I_{t}^{V}(\mu,\phi)\|_{\mathbb{B}}&\leq\frac{c_{1}}{\sqrt{t}}\int_{\mathbb{R}^{d}}{\alpha\big{(}t^{\frac{1}{2}}\big{)}|\phi(x)|\mu(\text{\rm{d}}x)}\\ &{\leq\frac{c_{1}}{\sqrt{t}}\alpha\big{(}t^{\frac{1}{2}}\big{)}\|\phi\|_{L^{k}(\mu)},\ \ t\in(0,T],\mu\in\mathscr{P}_{k},\phi\in T_{\mu,k}}.\end{split}$
To estimate $\tilde{I}_{t}^{V}(\mu,v,\tilde{v})$, noting that $X_{t}^{\mu}$ solves (3.9) for $\gamma=\mu$ and initial value $X_{0}^{\mu}$ replacing $x$, by the standard Markov property for solutions of (3.9), we have
$\tilde{I}_{t}^{V,1}(\mu,v,\tilde{v})=\int_{\mathbb{R}^{d}}\mathbb{E}\bigg{[}V(X_{t}^{x,\mu})\int_{0}^{t}{\frac{\alpha(s^{\frac{1}{2}})}{\{\tilde{\alpha}(s^{\frac{1}{2}})s\}^{\frac{1}{2}}}}$
×$\qquad\qquad\times\Big{\langle}\zeta_{s}(X_{s}^{x,\mu})\nabla_{v_{s}}^{\mathbb{B}}B_{s}(X_{s}^{x,\mu},P_{s}^{*}\mu,\cdot)(P_{s}V(\mu)),\ \text{\rm{d}}W_{s}\Big{\rangle}\bigg{]}\mu(\text{\rm{d}}x),$
$\tilde{I}_{t}^{V,2}(\mu,v,\tilde{v})=\int_{\mathbb{R}^{d}}\mathbb{E}\bigg{[}V(X_{t}^{x,\mu})\int_{0}^{t}\Big{\langle}\zeta_{s}(X_{s}^{x,\mu})$
×$\qquad\qquad\times\mathbb{E}\big{[}\big{\langle}\{D^{L}B_{s}(y,\cdot,P_{s}V(\mu))\}(P_{s}^{\ast}\mu)(X_{s}^{\mu}),(\nabla\Theta_{s}(X_{s}^{\mu}))^{-1}\tilde{v}_{s}\big{\rangle}\big{]}|_{y=X_{s}^{x,\mu}},\ \text{\rm{d}}W_{s}\Big{\rangle}\bigg{]}\mu(\text{\rm{d}}x).$
So, by $(B_{2})$, $(B_{3})$, (2.22) in Lemma 2.5 for $p=2$ and $y=\tilde{\theta}^{\lambda_{0},\mu,B}_{t}(x)$, (2.1), (3.24) and Jensen’s inequality, we find a constant $c_{2}>0$ such that
(3.27) $\begin{split}&\|\tilde{I}_{t}^{V,1}(\mu,v,\tilde{v})\|_{\mathbb{B}}\leq c_{2}{\alpha\big{(}t^{\frac{1}{2}}\big{)}}\bigg{(}\int_{0}^{t}{\frac{\alpha(s^{\frac{1}{2}})^{2}}{s\tilde{\alpha}(s^{\frac{1}{2}})}}\|v_{s}\|^{2}_{\mathbb{B}}\text{\rm{d}}s\bigg{)}^{\frac{1}{2}},\\ &\|\tilde{I}_{t}^{V,2}(\mu,v,\tilde{v})\|_{\mathbb{B}}\leq c_{2}{\alpha\big{(}t^{\frac{1}{2}}\big{)}}\left(\int_{0}^{t}(\mathbb{E}|\tilde{v}_{s}|^{k})^{\frac{2}{k}}\text{\rm{d}}s\right)^{\frac{1}{2}}.\end{split}$
Combining this with (3.19), (3.20) and (3.26), we find a constant $c_{3}>0$ such that
(3.28) $\begin{split}\|\bar{H}_{t}(v,\tilde{v})\|_{\mathbb{B}}\leq&\,c_{3}\|\phi\|_{L^{k}(\mu)}\sqrt{\tilde{\alpha}\big{(}t^{\frac{1}{2}}\big{)}}\\ &+c_{3}\sqrt{t\tilde{\alpha}\big{(}t^{\frac{1}{2}}\big{)}}\bigg{(}\int_{0}^{t}\frac{\alpha(s^{\frac{1}{2}})^{2}}{s\tilde{\alpha}\big{(}s^{\frac{1}{2}}\big{)}}\|v_{s}\|^{2}_{\mathbb{B}}\text{\rm{d}}s\bigg{)}^{\frac{1}{2}}\\ &+c_{3}\sqrt{t\tilde{\alpha}\big{(}t^{\frac{1}{2}}\big{)}}\left(\int_{0}^{t}(\mathbb{E}|\tilde{v}_{s}|^{k})^{\frac{2}{k}}\text{\rm{d}}s\right)^{\frac{1}{2}}.\end{split}$
Then (3.22) holds.
(b) We intend to prove that $\Phi^{(n)}$ in (i) has a unique fixed point in $\mathscr{V}_{0}\times\mathscr{C}_{k,n}.$ Obviously, for any $\delta>0$, $\mathscr{V}_{0}\times\mathscr{C}_{k}$ is complete under the metric
$\rho_{\delta,n}\big{(}(v,\tilde{v}),(w,\tilde{w})\big{)}:=\sup_{t\in[0,T]}\text{\rm{e}}^{-\delta t}\|v_{t}-w_{t}\|_{\mathbb{B}}+\rho_{\delta,n}^{(k)}\big{(}\tilde{v},\tilde{w}\big{)},$
$\rho_{\delta,n}^{(k)}\big{(}\tilde{v},\tilde{w}\big{)}:=\bigg{(}\mathbb{E}\Big{[}\sup_{t\in[0,\tau_{n}\land T]}\text{\rm{e}}^{-\delta kt}|\tilde{v}_{t}-\tilde{w}_{t}|^{k}\Big{]}\bigg{)}^{\frac{1}{k}},\ \ v,w\in\mathscr{V}_{0},\tilde{v},\tilde{w}\in\mathscr{C}_{k}.$
So, it suffices to prove the contraction of $\Phi^{(n)}$ in $\rho_{\delta,n}$ for large enough $\delta>0$.
By the definition of $H_{t}$, $\|\nabla(B+\lambda_{0}u)\|_{\infty}<\infty$, and repeating step (2) in the proof of [29, Lemma 4.1], we find a function $c_{n}:(0,\infty)\rightarrow(0,\infty)$ with $c_{n}(\delta)\rightarrow 0$ as $\delta\rightarrow\infty$, such that
(3.29) $\begin{split}&\rho_{\delta,n}^{(k)}(H(v,\tilde{v}),H(w,\tilde{w})):=\bigg{(}\mathbb{E}\Big{[}\sup_{t\in[0,\tau_{n}]}\text{\rm{e}}^{-\delta kt}|H_{t}(v,\tilde{v})-H_{t}(w,\tilde{w})|^{k}\Big{]}\bigg{)}^{\frac{1}{k}}\\ &\leq\sup_{t\in[0,T]}\text{\rm{e}}^{-\delta t}\|v_{t}-w_{t}\|_{\mathbb{B}}\sup_{s\in[0,T]}\int_{0}^{s}\text{\rm{e}}^{-\delta(s-t)}{\frac{\alpha(t^{\frac{1}{2}})}{\{\tilde{\alpha}(t^{\frac{1}{2}})t\}^{\frac{1}{2}}}}\text{\rm{d}}t\\ &+c(\delta,n)\rho_{\delta,n}^{k}(\tilde{v},\tilde{w}),\ \ n\geq 1,(v,\tilde{v}),(w,\tilde{w})\in\mathscr{V}_{0}\times\mathscr{C}_{k,n},\ \delta>0.\end{split}$
Moreover, by (3.20), (3.26) and (3.27), we find a constant $c_{4}>0$ such that
$\|\bar{H}_{t}(v,\tilde{v})-\bar{H}_{t}(w,\tilde{w})\|_{\mathbb{B}}=\frac{\{t\tilde{\alpha}\big{(}t^{\frac{1}{2}}\big{)}\}^{\frac{1}{2}}}{\alpha\big{(}t^{\frac{1}{2}}\big{)}}\|\tilde{I}_{t}^{V}(\mu,v-w,\tilde{v}-\tilde{w})\|_{\mathbb{B}}$
$\leq c_{4}\bigg{(}\int_{0}^{t}\frac{\alpha(s^{\frac{1}{2}})^{2}}{s\tilde{\alpha}(s^{\frac{1}{2}})}\|v_{s}-w_{s}\|^{2}_{\mathbb{B}}\text{\rm{d}}s\bigg{)}^{\frac{1}{2}}+c_{4}\left(\int_{0}^{t}(\mathbb{E}|\tilde{v}_{s}-\tilde{w}_{s}|^{k})^{\frac{2}{k}}\text{\rm{d}}s\right)^{\frac{1}{2}},$
$\qquad\ v,w\in\mathscr{V}_{0},\tilde{v},\tilde{w}\in\mathscr{C}_{k,n},t\in[0,T].$
Combining this with (3.19) and (3.29), we conclude that $\Phi^{(n)}$ is contractive in the complete metric space $(\mathscr{V}_{0}\times\mathscr{C}_{k,n},\rho_{\delta,n})$ for large enough $\delta>0$, and hence has a unique fixed point.
(2) Proof of (ii). Combining the first equation in (3.14), (B), Itô’s formula, the maximal inequality in [30, Lemma 2.1], Krylov’s estimate (see [32, Theorem 3.1]) and hence Khasminskii’s estimate (see [30, Lemma 4.1]), for any $j\geq k$ we find a constant $c_{1}>0$ such that
d $|{\tilde{v}}^{n}_{t}|^{j+1}$
$\leq c_{1}\bigg{\{}|{\tilde{v}}^{n}_{t}|^{j+1}\text{\rm{d}}\mathscr{L}_{t}+\big{(}\mathbb{E}[|\tilde{v}_{t}^{n}|^{k}]\big{)}^{\frac{j+1}{k}}\text{\rm{d}}t$
$\qquad\quad+|{\tilde{v}}_{t}^{n}|^{j}{\frac{\alpha(t^{\frac{1}{2}})}{\{t\tilde{\alpha}(t^{\frac{1}{2}})\}^{\frac{1}{2}}}}\|v_{t}^{(n)}\|_{\mathbb{B}}\text{\rm{d}}t+\text{\rm{d}}M_{t}\bigg{\}},\ \ t\in[0,\tau_{n}]$
for some local martingale $M_{t}$ and an adapted increasing process $\mathscr{L}_{t}$ with $\mathbb{E}[\text{\rm{e}}^{p\mathscr{L}_{T}}]<\infty$ for any $p>0$. Then by the stochastic Gronwall inequality (see [31, Lemma 3.7]), for any $j\geq k$, we find constants $c_{2},c_{3}>0$ such that
$\mathbb{E}\Big{[}\sup_{s\in[0,t]}|{\tilde{v}}_{s}^{(n)}|^{j}\Big{|}\mathscr{F}_{0}\Big{]}\leq c_{2}|\phi(X_{0}^{\mu})|^{j}+c_{2}\bigg{(}\int_{0}^{t}\big{(}\mathbb{E}[|\tilde{v}_{s}^{(n)}|^{k}]\big{)}^{\frac{j+1}{k}}\text{\rm{d}}s\bigg{)}^{\frac{j}{j+1}}$
$\qquad\qquad\quad+c_{2}\left(\mathbb{E}\bigg{(}\int_{0}^{t}|{\tilde{v}}_{t}^{n}|^{j}{\frac{\alpha(s^{\frac{1}{2}})}{(\tilde{\alpha}(s^{\frac{1}{2}})s)^{\frac{1}{2}}}}\|v_{s}^{(n)}\|_{\mathbb{B}}\text{\rm{d}}s\bigg{|}\mathscr{F}_{0}\bigg{)}\right)^{\frac{j}{j+1}}$
$\qquad\qquad\quad\leq c_{3}|\phi(X_{0}^{\mu})|^{j}+c_{3}\bigg{(}\int_{0}^{t}\big{(}\mathbb{E}[|\tilde{v}_{s}^{(n)}|^{k}]\big{)}^{\frac{j+1}{k}}\text{\rm{d}}s\bigg{)}^{\frac{j}{j+1}}$
$\qquad\qquad\quad+c_{3}\left(\int_{0}^{t}{\frac{\alpha(s^{\frac{1}{2}})}{(\tilde{\alpha}(s^{\frac{1}{2}})s)^{\frac{1}{2}}}}\|v_{s}^{(n)}\|_{\mathbb{B}}\text{\rm{d}}s\right)^{j}+\frac{1}{2}\mathbb{E}\Big{[}\sup_{s\in[0,t]}|{\tilde{v}}_{s}^{(n)}|^{j}\Big{|}\mathscr{F}_{0}\Big{]}.$
Noting that (3.19) and the Schwartz inequality imply
$\bigg{(}\int_{0}^{t}{\frac{\alpha(s^{\frac{1}{2}})}{(\tilde{\alpha}(s^{\frac{1}{2}})s)^{\frac{1}{2}}}}\|v_{s}^{(n)}\|_{\mathbb{B}}\text{\rm{d}}s\bigg{)}^{2}\leq 4{\tilde{\alpha}\big{(}t^{\frac{1}{2}}\big{)}}\int_{0}^{t}\|v_{s}^{(n)}\|_{\mathbb{B}}^{2}\text{\rm{d}}s,$
we find a constant $c>0$ such that
(3.30) $\begin{split}&\mathbb{E}\Big{[}\sup_{s\in[0,t]}|{\tilde{v}}_{s}^{(n)}|^{j}\Big{|}\mathscr{F}_{0}\Big{]}\leq c|\phi(X_{0}^{\mu})|^{j}+c\bigg{(}\int_{0}^{t}\big{(}\mathbb{E}[|\tilde{v}_{s}^{(n)}|^{k}]\big{)}^{\frac{j+1}{k}}\text{\rm{d}}s\bigg{)}^{\frac{j}{j+1}}\\ &\qquad+c\left({\tilde{\alpha}\big{(}t^{\frac{1}{2}}\big{)}}\int_{0}^{t}\|v_{s}^{(n)}\|_{\mathbb{B}}^{2}\text{\rm{d}}s\right)^{\frac{j}{2}},\ \ t\in[0,T].\end{split}$
Choosing $j=k$, taking expectations to both sides, and noting that
(3.31) $\begin{split}&c\bigg{(}\int_{0}^{t}\big{(}\mathbb{E}[|\tilde{v}_{s}^{(n)}|^{k}]\big{)}^{\frac{k+1}{k}}\text{\rm{d}}s\bigg{)}^{\frac{k}{k+1}}\\ &\leq c_{4}\bigg{(}\int_{0}^{t}\big{(}\mathbb{E}[|\tilde{v}_{s}^{(n)}|^{k}]\big{)}^{2}\text{\rm{d}}s\bigg{)}^{\frac{1}{2}}\leq\frac{1}{2}\mathbb{E}\Big{[}\sup_{s\in[0,t]}|{\tilde{v}}_{s}^{(n)}|^{k}\Big{]}+\frac{c_{4}^{2}}{2}\int_{0}^{t}\mathbb{E}[|\tilde{v}_{s}^{(n)}|^{k}]\text{\rm{d}}s\end{split}$
holds for some constant $c_{4}>c$, we derive
$\mathbb{E}\Big{[}\sup_{s\in[0,t]}|{\tilde{v}}_{s}^{(n)}|^{k}\Big{]}\leq$  $\,2c_{4}\|\phi\|_{L^{k}(\mu)}^{k}+c_{4}^{2}\int_{0}^{t}\mathbb{E}[|\tilde{v}_{s}^{(n)}|^{k}]\text{\rm{d}}s$
$+2c_{4}\left({\tilde{\alpha}\big{(}t^{\frac{1}{2}}\big{)}}\int_{0}^{t}\|v_{s}^{(n)}\|_{\mathbb{B}}^{2}\text{\rm{d}}s\right)^{\frac{k}{2}}.$
So, Gronwall’s inequality yields
(3.32) $\begin{split}&\mathbb{E}\Big{[}\sup_{s\in[0,t]}|{\tilde{v}}_{s}^{(n)}|^{k}\Big{]}\leq 2c_{4}\text{\rm{e}}^{c_{4}^{2}T}\bigg{\{}\|\phi\|_{L^{k}(\mu)}^{k}+\left({\tilde{\alpha}\big{(}t^{\frac{1}{2}}\big{)}}\int_{0}^{t}\|v_{s}^{(n)}\|_{\mathbb{B}}^{2}\text{\rm{d}}s\right)^{\frac{k}{2}}\bigg{\}}.\end{split}$
On the other hand, by taking $(v,\tilde{v})=(v^{(n)},\tilde{v}^{(n)})$ in (3.28) and noting that $\bar{H}_{t}(v,\tilde{v})=v_{t}^{(n)}$ due to (3.21), we find a constant $c_{5}>0$ such that
$\|v_{t}^{(n)}\|_{\mathbb{B}}^{2}\leq c_{5}\tilde{\alpha}\big{(}t^{\frac{1}{2}}\big{)}\|\phi\|_{L^{k}(\mu)}^{2}+c_{5}\int_{0}^{t}\frac{\alpha(s^{\frac{1}{2}})^{2}}{s\tilde{\alpha}(s^{\frac{1}{2}})}\|v_{s}^{(n)}\|^{2}_{\mathbb{B}}\text{\rm{d}}s+c_{5}\int_{0}^{t}(\mathbb{E}|\tilde{v}_{s}^{(n)}|^{k})^{\frac{2}{k}}\text{\rm{d}}s,\ \ t\in[0,T].$
Combining this with (3.19), (3.32) and Gronwall’s inequality, we find a constant $c_{6}>0$ such that for any $n\geq 1$ and $t\in[0,T]$,
$\|v_{t}^{(n)}\|_{\mathbb{B}}\leq c_{6}\|\phi\|_{L^{k}(\mu)}\,\sqrt{\tilde{\alpha}(t^{\frac{1}{2}})},$
$\mathbb{E}\Big{[}\sup_{s\in[0,t\land\tau_{n}]}|{\tilde{v}}_{s}^{(n)}|^{k}\Big{]}\leq c_{6}\|\phi\|_{L^{k}(\mu)}^{k},\ \ \mu\in\mathscr{P}_{k},\ \phi\in T_{\mu,k}.$
This together with (3.30) proves (ii). ∎
3.3 Proof of Theorem 3.2
By Lemma 3.4, the proof of Theorem 3.2(1)-(2) is completed by the following lemma.
Lemma 3.5.
[…]
$\sup_{\|\phi\|_{L^{k}(\mu)}\leq 1}\bigg{\{}\sup_{t\in(0,T]}\frac{\sqrt{t}}{\alpha(t^{\frac{1}{2}})}\|D_{\phi}P_{t}V(\mu)\|_{\mathbb{B}}+\mathbb{E}\Big{[}\sup_{t\in[0,T]}|\nabla_{\phi}X_{t}^{\mu}|^{k}\Big{]}\bigg{\}}\leq c.$
Proof.
By Lemma 3.4, it suffices to prove (3.15). For any $t\in[0,T]$ and $\varepsilon\in(0,1]$, (3.10) for $V$ replacing $f$ implies
(3.33) $\begin{split}&v_{t}^{\varepsilon}:=\frac{P_{t}V(\mu_{\varepsilon})-P_{t}V(\mu)}{\varepsilon}=v_{t}^{\varepsilon,1}+v_{t}^{\varepsilon,2},\\ &v_{t}^{\varepsilon,1}:=\frac{P_{t}^{\mu}V(\mu_{\varepsilon})-P_{t}^{\mu}V(\mu)}{\varepsilon},\ \ v_{t}^{\varepsilon,2}:=\frac{P_{t}^{\mu_{\varepsilon}}V(\mu_{\varepsilon})-P_{t}^{\mu}V(\mu_{\varepsilon})}{\varepsilon}.\end{split}$
Next, let
(3.34) $\begin{split}&\hat{v}_{t}:={\frac{\alpha(t^{\frac{1}{2}})}{\{\tilde{\alpha}(t^{\frac{1}{2}})t\}^{\frac{1}{2}}}}v_{t}(\mu,\phi),\\ &\tilde{v}_{t}:=\tilde{v}_{t}(\mu,\phi),\ \ \tilde{v}_{t}^{\varepsilon}=\frac{\Theta_{t}(X_{t}^{\mu_{\varepsilon}})-\Theta_{t}(X_{t}^{\mu})}{\varepsilon}.\end{split}$
Then (3.15) follows from
(3.35) $\lim_{\varepsilon\rightarrow 0}\sup_{t\in(0,T]}\frac{\sqrt{t}}{\alpha(t^{\frac{1}{2}})}\big{\|}v_{t}^{\varepsilon}-\hat{v}_{t}\big{\|}_{\mathbb{B}}=0,\ \ \lim_{\varepsilon\rightarrow 0}\mathbb{E}\Big{[}\sup_{t\in[0,T]}|\tilde{v}_{t}^{\varepsilon}-\tilde{v}_{t}|^{k}\Big{]}=0.$
In the following, we first estimate $\|v_{s}^{\varepsilon,1}\|_{\mathbb{B}}$, $\|v_{s}^{\varepsilon,2}\|_{\mathbb{B}}$, $|\tilde{v}_{s}^{\varepsilon}|$, $\|v_{t}^{\varepsilon}-\hat{v}_{t}\big{\|}_{\mathbb{B}}$ and $|\tilde{v}_{t}^{\varepsilon}-\tilde{v}_{t}|$ in (a)-(e), then we prove (3.35) in (f).
(a) For $v_{t}^{\varepsilon,1}$. Let
$V_{N}:=\frac{(N\land\|V\|_{\mathbb{B}})V}{\|V\|_{\mathbb{B}}},\ \ N\geq 1.$
By (3.11), we have
$\lim_{\varepsilon\downarrow 0}\frac{P_{t}^{\mu}V_{N}(\mu_{\varepsilon})-P_{t}^{\mu}V_{N}(\mu)}{\varepsilon}=I_{t}^{V_{N}}(\mu,\phi),$
$P_{t}^{\mu}V_{N}(\mu_{\varepsilon})-P_{t}^{\mu}V_{N}(\mu)=\int_{0}^{\varepsilon}\frac{\text{\rm{d}}}{\text{\rm{d}}r}P_{t}^{\mu}V_{N}(\mu_{r})\text{\rm{d}}r=\int_{0}^{\varepsilon}I_{t}^{V_{N}}(r,\mu,\phi)\text{\rm{d}}r,\ \ N\geq 1.$
Since $(B_{2})$ implies
(3.36) ${\|V_{N}\|_{\mathbb{B}}\leq\|V\|_{\mathbb{B}}\leq c(1+|\cdot|^{k-1})}$
for some constant $c>0$, by (2.14), (3.24), the same argument to (3.26), and the dominated convergence theorem for $N\rightarrow\infty$, we find a constant $c(\mu,\phi)>0$ depending on $\phi,\mu$ such that
(3.37) $\|v_{t}^{\varepsilon,1}\|_{\mathbb{B}}=\frac{\|P_{t}^{\mu}V(\mu_{\varepsilon})-P_{t}^{\mu}V(\mu)\|_{\mathbb{B}}}{\varepsilon}\leq\frac{c(\mu,\phi)\alpha(t^{\frac{1}{2}})}{\sqrt{t}},\ \ \ \varepsilon\in(0,1],t\in(0,T],$
and
$\lim_{\varepsilon\downarrow 0}\Big{\|}\frac{P_{t}^{\mu}V(\mu_{\varepsilon})-P_{t}^{\mu}V(\mu)}{\varepsilon}-I_{t}^{V}(\mu,\phi)\Big{\|}_{\mathbb{B}}$
$\leq\lim_{N\rightarrow\infty}\lim_{\varepsilon\downarrow 0}\bigg{\{}\|I_{t}^{V-V_{N}}(\mu,\phi)\|_{\mathbb{B}}+\frac{\|P_{t}^{\mu}(V-V_{N})(\mu_{\varepsilon})-P_{t}^{\mu}(V-V_{N})(\mu)\|_{\mathbb{B}}}{\varepsilon}\bigg{\}}$
$\leq\lim_{N\rightarrow\infty}\frac{c\|\phi\|_{L^{k}(\mu)}}{\sqrt{t}}\bigg{(}\int_{\mathbb{R}^{d}}\mathbb{E}[\|(V-V_{N})(X_{t}^{x,\mu})\|_{\mathbb{B}}^{k^{*}}]\mu(\text{\rm{d}}x)\bigg{)}^{\frac{1}{k^{*}}}=0,\ \ t\in(0,T].$
Hence,
$h_{t,1}(\varepsilon):=\big{\{}v_{t}^{\varepsilon,1}-I_{t}^{V}(\mu,\phi)\big{\}}\frac{\{t\tilde{\alpha}(t^{\frac{1}{2}})\}^{\frac{1}{2}}}{\alpha(t^{\frac{1}{2}})}$
satisfies
(3.38) $\lim_{\varepsilon\rightarrow 0}\|h_{t,1}(\varepsilon)\|_{\mathbb{B}}=0,\ \ \sup_{\varepsilon\in(0,1]}\|h_{t,1}(\varepsilon)\|_{\mathbb{B}}\leq 2c(\mu,\phi)\sqrt{\tilde{\alpha}(t^{\frac{1}{2}})},\ \ t\in[0,T].$
(b) For $v_{t}^{\varepsilon,2}$ and $v_{t}^{\varepsilon}$. Recall that $X_{t}^{\mu_{\varepsilon},\mu}$ solves (3.2) with initial value $X_{0}^{\varepsilon}$. Let
(3.39) $\begin{split}&R_{t}^{\varepsilon}:=\text{\rm{e}}^{\int_{0}^{t}\langle\eta_{s}^{\varepsilon},\text{\rm{d}}W_{s}\rangle-\frac{1}{2}\int_{0}^{t}|\eta_{s}^{\varepsilon}|^{2}\text{\rm{d}}s},\\ &\eta_{s}^{\varepsilon}:=\zeta_{s}(X_{s}^{\mu_{\varepsilon},\mu})\\ &\qquad\quad\times\big{\{}B_{s}(X_{s}^{\mu_{\varepsilon},\mu},P_{s}^{\ast}\mu_{\varepsilon},P_{s}V(\mu_{\varepsilon}))-B_{s}(X_{s}^{\mu_{\varepsilon},\mu},P_{s}^{\ast}\mu,P_{s}V(\mu))\big{\}},\ \ s,t\in[0,T].\end{split}$
Since (B) implies (A), (2.24) holds so that there exists a constant $c>0$ such that
$\sup_{x\in\mathbb{R}^{d},z\in\mathbb{B},t\in[0,T]}|B_{t}(x,P_{s}^{*}\mu_{\varepsilon},z)-B_{t}(x,P_{s}^{*}\mu,z)|\leq c\mathbb{W}_{k}(\mu_{\varepsilon},\mu)\leq c\varepsilon\|\phi\|_{L^{k}(\mu)},\ \ \varepsilon\in[0,1].$
Hence, by (B) we find a constant $c(\mu,\phi)>0$ such that
(3.40) $\begin{split}|\eta_{s}^{\varepsilon}|&\leq K\|P_{s}V(\mu_{\varepsilon})-P_{s}V(\mu)\|_{\mathbb{B}}+c\varepsilon\|\phi\|_{L^{k}(\mu)}\\ &\leq c(\mu,\phi)\varepsilon\big{(}1+\|v_{s}^{\varepsilon}\|_{\mathbb{B}}\big{)},\ \ s\in[0,T],\ \varepsilon\in(0,1].\end{split}$
Then by Girsanov’s theorem,
$W_{t}^{\varepsilon}:=W_{t}-\int_{0}^{t}\eta_{s}^{\varepsilon}\text{\rm{d}}s,\ \ s\in[0,T]$
is a Brownian motion under $\mathbb{Q}:=R_{T}^{\varepsilon}\mathbb{P}$. Reformulate (3.9) with $\gamma=\mu$ as
$\text{\rm{d}}X_{t}^{\mu_{\varepsilon},\mu}$ =$=\Big{\{}b_{t}^{(0)}(X_{t}^{\mu_{\varepsilon},\mu})+B_{t}(X_{t}^{\mu_{\varepsilon},\mu},P_{t}^{*}\mu,P_{t}V(\mu))\Big{\}}\text{\rm{d}}t+\sigma_{t}(X_{t}^{\mu_{\varepsilon},\mu})\text{\rm{d}}W_{t}$
=$=\Big{\{}b_{t}^{(0)}(X_{t}^{\mu_{\varepsilon},\mu})+B_{t}(X_{t}^{\mu_{\varepsilon},\mu},P_{t}^{*}\mu_{\varepsilon},P_{t}V(\mu_{\varepsilon}))\Big{\}}\text{\rm{d}}t+\sigma_{t}(X_{t}^{\mu_{\varepsilon},\mu})\text{\rm{d}}W_{t}^{\varepsilon},$
$\qquad\ X_{0}^{\mu_{\varepsilon},\mu}=X_{0}^{\mu_{\varepsilon},\mu_{\varepsilon}}=X_{0}^{\mu_{\varepsilon}}=X_{0}^{\mu}+\varepsilon\phi(X_{0}^{\mu}).$
By the weak uniqueness of (3.9) with $\gamma=\mu_{\varepsilon},$ we get
(3.41) $v_{t}^{\varepsilon,2}:=\frac{P_{t}^{\mu_{\varepsilon}}V(\mu_{\varepsilon})-P_{t}^{\mu}V(\mu_{\varepsilon})}{\varepsilon}=\frac{1}{\varepsilon}\mathbb{E}[V(X_{t}^{\mu_{\varepsilon},\mu})(R_{t}^{\varepsilon}-1)],\ \ t\in[0,T].$
Combinig this with $[V]_{\alpha}\leq 1,$ and using Lemma 2.5 for $p=k^{\ast}$, we obtain
(3.42) $\begin{split}&\Big{\|}v_{t}^{\varepsilon,2}-\frac{1}{\varepsilon}\mathbb{E}[V(X_{t}^{\mu})(R_{t}^{\varepsilon}-1)]\Big{\|}_{\mathbb{B}}\leq\frac{1}{\varepsilon}\mathbb{E}[\alpha(|X_{t}^{\mu}-X_{t}^{\mu_{\varepsilon},\mu}|)|R_{t}^{\varepsilon}-1|]\\ &\leq\frac{1}{\varepsilon}\Big{(}\mathbb{E}[|R_{t}^{\varepsilon}-1|^{k^{*}}]\Big{)}^{\frac{1}{k^{*}}}\alpha\Big{(}\big{(}\mathbb{E}[|X_{t}^{\mu}-X_{t}^{\mu_{\varepsilon},\mu}|^{k}]\big{)}^{\frac{1}{k}}\Big{)}.\end{split}$
By (3.24) and (3.36), we find a constant $c_{0}(\mu,\phi)>0$ such that
(3.43) $\mathbb{E}[\|V(X_{t}^{\mu_{\varepsilon},\mu})\|_{\mathbb{B}}^{k}]\leq c_{0}(\mu,\phi),\ \ t\in[0,T],\varepsilon\in(0,1].$
On the other hand, by (3.40), for any $p\geq 1$ there exists a constant $c(p,\mu,\phi)>0$ such that
(3.44) $\mathbb{E}[|R_{t}^{\varepsilon}-1|^{p}]\leq c(p,\mu,\phi)\varepsilon^{p}\bigg{(}\int_{0}^{t}\big{(}1+\|v_{s}^{\varepsilon}\|_{\mathbb{B}}^{2}\big{)}\text{\rm{d}}s\bigg{)}^{\frac{p}{2}},\ \ t\in[0,T],\varepsilon\in[0,1].$
Combining this with (3.41) and (3.43), we find a constant $c_{1}(\mu,\phi)>0$ such that
$\|v_{t}^{\varepsilon,2}\|_{\mathbb{B}}$ $\leq\big{(}\mathbb{E}[\|V(X_{t}^{\mu_{\varepsilon},\mu})\|_{\mathbb{B}}^{k}]\big{)}^{\frac{1}{k}}\frac{\big{(}\mathbb{E}[|R_{t}^{\varepsilon}-1|^{k^{*}}]\big{)}^{\frac{1}{k^{*}}}}{\varepsilon}$
$\leq c_{1}(\mu,\phi)\bigg{(}\int_{0}^{t}\big{(}1+\|v_{s}^{\varepsilon}\|_{\mathbb{B}}^{2}\big{)}\text{\rm{d}}s\bigg{)}^{\frac{1}{2}}.$
This together with (3.33) and (3.37) yields that for some constant $c_{2}(\mu,\phi)>0$,
$\|v_{t}^{\varepsilon}\|_{\mathbb{B}}^{2}\leq 2\|v_{t}^{\varepsilon,1}\|_{\mathbb{B}}^{2}+2\|v_{t}^{\varepsilon,2}\|_{\mathbb{B}}^{2}\leq c_{2}(\mu,\phi)\bigg{(}\frac{\alpha(t^{\frac{1}{2}})^{2}}{t}+\int_{0}^{t}\|v_{s}^{\varepsilon}\|_{\mathbb{B}}^{2}\text{\rm{d}}s\bigg{)},\ \ t\in[0,T].$
By Grownwall’s inequality, we find a constant $c_{3}(\mu,\phi)>0$ such that
(3.45) $\begin{split}&\|v_{t}^{\varepsilon}\|_{\mathbb{B}}^{2}\leq c_{2}(\mu,\phi)\bigg{(}\frac{\alpha(t^{\frac{1}{2}})^{2}}{t}+\int_{0}^{t}\frac{\alpha(s^{\frac{1}{2}})^{2}}{s}\text{\rm{e}}^{c_{2}(\mu,\phi)(t-s)}\text{\rm{d}}s\bigg{)}\\ &\leq c_{3}(\mu,\phi)\frac{\alpha(t^{\frac{1}{2}})^{2}}{t},\ \ t\in[0,T],\ \ \varepsilon\in(0,1].\end{split}$
(c) For $|\tilde{v}_{s}^{\varepsilon}|$. Recall that $\Theta_{t}=u_{t}+id$ for $u_{t}=u_{t}^{\lambda_{0},\mu}$. Similarly to (2.16) and (2.21), by Itô’s formula for $\Theta_{t}(X_{t}^{\mu_{\varepsilon}})-\Theta_{t}(X_{t}^{\mu})$, we obtain
(3.46) $\begin{split}&\text{\rm{d}}\tilde{v}_{t}^{\varepsilon}=\frac{1}{\varepsilon}\Big{\{}\big{\{}(\nabla\Theta_{t})\sigma_{t}\big{\}}(X_{t}^{\mu_{\varepsilon}})-\big{\{}(\nabla\Theta_{t})\sigma_{t}\big{\}}(X_{t}^{\mu})\Big{\}}\text{\rm{d}}W_{t}\\ &+\frac{1}{\varepsilon}\Big{[}\lambda_{0}u_{t}(X_{t}^{\mu_{\varepsilon}})-\lambda_{0}u_{t}(X_{0}^{\mu})+\nabla_{B_{t}(X_{t}^{\mu_{\varepsilon}},P_{t}^{*}\mu_{\varepsilon},P_{t}V(\mu_{\varepsilon}))-B_{t}(X_{t}^{\mu_{\varepsilon}},P_{t}^{*}\mu,P_{t}V(\mu))}\Theta_{t}(X_{t}^{\mu_{\varepsilon}})\Big{]}\text{\rm{d}}t\\ &+\frac{1}{\varepsilon}[B_{t}(X_{t}^{\mu_{\varepsilon}},P_{t}^{*}\mu,P_{t}V(\mu))-B_{t}(X_{t}^{\mu},P_{t}^{*}\mu,P_{t}V(\mu))]\text{\rm{d}}t,\\ &\tilde{v}_{0}^{\varepsilon}=\frac{\Theta_{0}(X_{0}^{\mu_{\varepsilon}})-\Theta_{0}(X_{0}^{\mu})}{\varepsilon}.\end{split}$
Similarly to (3.30), by combining this with (B), (2.9), (3.45) and the stochastic Gronwall inequality, for any $j\geq 1$ we find constants $c(j),c_{4}(\mu,\phi)>0$ such that
(3.47) $\begin{split}&\mathbb{E}\Big{[}\sup_{s\in[0,t]}|\tilde{v}_{s}^{\varepsilon}|^{j}\Big{|}\mathscr{F}_{0}\Big{]}\leq c(j)|\phi(X_{0}^{\mu})|^{j}+c_{4}(\mu,\phi)\\ &\qquad+c(j)\bigg{(}\int_{0}^{t}\big{(}\mathbb{E}[|\tilde{v}_{s}^{\varepsilon}|^{k}]\big{)}^{\frac{j+1}{k}}\text{\rm{d}}s\bigg{)}^{\frac{j}{j+1}},\ \ t\in[0,T],\varepsilon\in(0,1].\end{split}$
Letting $j=k$, taking expectation to both sides and noting that
$c(j)\bigg{(}\int_{0}^{t}\big{(}\mathbb{E}[|\tilde{v}_{s}^{\varepsilon}|^{k}]\big{)}^{\frac{k+1}{k}}\text{\rm{d}}s\bigg{)}^{\frac{k}{k+1}}$
$\leq\tilde{c}(j)\bigg{(}\int_{0}^{t}\big{(}\mathbb{E}[|\tilde{v}_{s}^{\varepsilon}|^{k}]\big{)}^{2}\text{\rm{d}}s\bigg{)}^{\frac{1}{2}}\leq\frac{1}{2}\mathbb{E}\Big{[}\sup_{s\in[0,t]}|\tilde{v}_{s}^{\varepsilon}|^{k}\Big{]}+\frac{\tilde{c}(j)^{2}}{2}\int_{0}^{t}\mathbb{E}\Big{[}\sup_{s\in[0,t]}|\tilde{v}_{s}^{\varepsilon}|^{k}\Big{]}\text{\rm{d}}s,$
for some $\tilde{c}(j)\geq c(j)$ as in (3.31), by Gronwall’s inequality we find a constant $c_{5}(\mu,\phi)>0$ such that
(3.48) $\mathbb{E}\Big{[}\sup_{s\in[0,T]}|\tilde{v}_{s}^{\varepsilon}|^{k}\Big{]}\leq c_{5}(\mu,\phi),\ \ \varepsilon\in(0,1],$
so that (3.47) implies
(3.49) $\mathbb{E}\Big{[}\sup_{s\in[0,T]}|\tilde{v}_{s}^{\varepsilon}|^{j}\Big{|}\mathscr{F}_{0}\Big{]}\leq c_{6}(\mu,\phi)+c(j)|\phi(X_{0}^{\mu})|^{j},\ \ \varepsilon\in(0,1]$
for some constant $c_{6}(\mu,\phi)>0.$
(d) For $|\tilde{v}_{s}^{\varepsilon}-\tilde{v}_{s}|^{k}$. Now, let
$\psi_{t}^{\varepsilon}:=\,$ $\mathbb{E}\Big{[}\big{\langle}\{D^{L}B_{t}(z,\cdot,P_{t}V(\mu))\}(P_{t}^{*}\mu)(X_{t}^{\mu}),(\nabla\Theta_{t}(X_{t}^{\mu}))^{-1}\tilde{v}_{t}^{\varepsilon}\big{\rangle}\Big{]}|_{z=X_{t}^{\mu}}$
+$\qquad+\big{\{}\nabla^{\mathbb{B}}_{v_{t}^{\varepsilon}}B_{t}(X_{t}^{\mu},P_{t}^{*}\mu,\cdot)\big{\}}(P_{t}V(\mu)).$
Then (3.46) can be reformulated as
$\tilde{v}_{t}^{\varepsilon}-\tilde{v}_{0}^{\varepsilon}$
$=\int_{0}^{t}\Big{\{}\nabla_{(\nabla\Theta_{s}(X_{s}^{\mu}))^{-1}\tilde{v}_{s}^{\varepsilon}}\big{[}B_{s}(\cdot,P_{s}^{*}\mu,P_{s}V(\mu))+\lambda_{0}u_{s}\big{]}(X_{s}^{\mu})+\nabla_{\psi_{s}^{\varepsilon}}\Theta_{s}(X_{s}^{\mu})\Big{\}}\text{\rm{d}}s$
$\quad+\int_{0}^{t}\big{[}\nabla_{(\nabla\Theta_{s}(X_{s}^{\mu}))^{-1}\tilde{v}_{s}^{\varepsilon}}\{(\nabla\Theta_{s})\sigma_{s}\}(X_{s}^{\mu})\big{]}\text{\rm{d}}W_{s}+\alpha_{t}^{\varepsilon},\ \ t\in[0,T],$
where for $t\in[0,T],$
$\alpha_{t}^{\varepsilon}:=\int_{0}^{t}\xi_{s}^{\varepsilon}\text{\rm{d}}s+\int_{0}^{t}\tilde{\xi}_{s}^{\varepsilon}\,\text{\rm{d}}W_{s},$
$\xi_{s}^{\varepsilon}:=\frac{1}{\varepsilon}\{[B_{s}(\cdot,P_{s}^{*}\mu_{\varepsilon},P_{s}V(\mu_{\varepsilon}))+\lambda_{0}u_{s}](X_{s}^{\mu_{\varepsilon}})-[B_{s}(\cdot,P_{s}^{*}\mu,P_{s}V(\mu))+\lambda_{0}u_{s}](X_{s}^{\mu})\}$
$\qquad\qquad+\nabla_{\frac{1}{\varepsilon}\{B_{s}(X_{s}^{\mu_{\varepsilon}},P_{s}^{*}\mu_{\varepsilon},P_{s}V(\mu_{\varepsilon}))-B_{s}(X_{s}^{\mu_{\varepsilon}},P_{s}^{*}\mu,P_{s}V(\mu))\}}u_{s}(X_{s}^{\mu_{\varepsilon}})$
$\qquad-\nabla_{(\nabla\Theta_{s}(X_{s}^{\mu}))^{-1}\tilde{v}_{s}^{\varepsilon}}\big{[}B_{s}(\cdot,P_{s}^{*}\mu,P_{s}V(\mu))+\lambda_{0}u_{s}\big{]}(X_{s}^{\mu})-\nabla_{\psi_{s}^{\varepsilon}}\Theta_{s}(X_{s}^{\mu}),$
$\tilde{\xi}_{s}^{\varepsilon}:=\frac{\{(\nabla\Theta_{s})\sigma_{s}\}(X_{s}^{\mu_{\varepsilon}})-\{(\nabla\Theta_{s})\sigma_{s}\}(X_{s}^{\mu})}{\varepsilon}-\nabla_{(\nabla\Theta_{s}(X_{s}^{\mu}))^{-1}\tilde{v}_{s}^{\varepsilon}}\{(\nabla\Theta_{s})\sigma_{s}\}(X_{s}^{\mu}).$
By the same argument leading to [29, (4.23),(4.26)] and (3.30) above, (B) together with (2.9), (3.45), (3.48), (3.49) implies
$\lim_{\varepsilon\rightarrow 0}\mathbb{E}\Big{[}\sup_{s\in[0,T]}|\alpha_{s}(\varepsilon)|^{j}\Big{|}\mathscr{F}_{0}\Big{]}=0,\ \ j\geq 1,$
and by Jensen’s inequality we find a constant $c>0$ such that
$\mathbb{E}\Big{[}\sup_{s\in[0,t]}|\tilde{v}_{s}^{\varepsilon}-\tilde{v}_{s}|^{k}\Big{|}\mathscr{F}_{0}\Big{]}\leq c|\tilde{v}_{0}^{\varepsilon}-\tilde{v}_{0}|^{k}+c\bigg{(}\mathbb{E}\Big{[}\sup_{s\in[0,t]}|\alpha_{s}(\varepsilon)|^{2k}\Big{|}\mathscr{F}_{0}\Big{]}\bigg{)}^{\frac{1}{2}}$
(3.50) $\quad+c\bigg{(}\int_{0}^{t}\|v_{s}^{\varepsilon}-\hat{v}_{s}\|_{\mathbb{B}}^{2}\text{\rm{d}}s\bigg{)}^{\frac{k}{2}}+c\bigg{(}\int_{0}^{t}\Big{\{}\mathbb{E}[|\tilde{v}_{s}^{\varepsilon}-\tilde{v}_{s}|^{k}]\Big{\}}^{2}\text{\rm{d}}s\bigg{)}^{\frac{1}{2}},\ \ t\in[0,T].$
(e) For $\|v_{t}^{\varepsilon}-\hat{v}_{t}\|_{\mathbb{B}}$. Similarly to (d), for $N_{s}$ and $\tilde{N}_{s}$ in (3.13), we have
(3.51) $\begin{split}&\frac{R_{t}^{\varepsilon}-1}{\varepsilon}=\int_{0}^{t}R_{s}^{\varepsilon}\big{\langle}\varepsilon^{-1}\eta_{s}^{\varepsilon},\text{\rm{d}}W_{s}\big{\rangle}\\ &=\int_{0}^{t}R_{s}^{\varepsilon}\Big{\langle}\frac{\zeta_{s}(X_{s}^{\mu_{\varepsilon},\mu})[B_{s}(X_{s}^{\mu_{\varepsilon},\mu},P_{s}^{*}\mu_{\varepsilon},P_{s}V(\mu_{\varepsilon}))-B_{s}(X_{s}^{\mu_{\varepsilon},\mu},P_{s}^{*}\mu,P_{s}V(\mu))]}{\varepsilon},\ \text{\rm{d}}W_{s}\Big{\rangle}\\ &=h_{t}(\varepsilon)+\int_{0}^{t}\Big{\langle}\zeta_{s}(X_{s}^{\mu})\big{\{}N_{s}(v^{\varepsilon})+\tilde{N}_{s}(\tilde{v}^{\varepsilon})\big{\}},\ \text{\rm{d}}W_{s}\Big{\rangle},\end{split}$
where
$h_{t}(\varepsilon):=\int_{0}^{t}\Big{\langle}$ $\zeta_{s}(X_{s}^{\mu})\Big{\{}R_{s}^{\varepsilon}\frac{B_{s}(X_{s}^{\mu_{\varepsilon},\mu},P_{s}^{*}\mu_{\varepsilon},P_{s}V(\mu_{\varepsilon}))-B_{s}(X_{s}^{\mu_{\varepsilon},\mu},P_{s}^{*}\mu,P_{s}V(\mu))}{\varepsilon}$
$\qquad\qquad-N_{s}(v^{\varepsilon})-\tilde{N}_{s}(\tilde{v}^{\varepsilon})\Big{\}},\ \text{\rm{d}}W_{s}\Big{\rangle}$
satisfies
(3.52) $\lim_{\varepsilon\rightarrow 0}\mathbb{E}\Big{[}\sup_{t\in[0,T]}|h_{t}(\varepsilon)|^{p}\Big{|}\mathscr{F}_{0}\Big{]}=0,\ \ p\geq 1.$
Noting that $X_{t}^{\mu}=X_{t}^{\mu,\mu}$, by (3.25) we find a constant $c(\mu,\phi)>0$ such that
(3.53) $\mathbb{E}\Big{[}\sup_{t\in[0,T]}|X_{t}^{\mu_{\varepsilon},\mu}-X_{t}^{\mu}|^{k}\Big{]}\leq c(\mu,\phi)\varepsilon^{k},\ \ \varepsilon\in[0,1],\mu\in\mathscr{P}_{k}.$
Combining these with (3.42), (3.45) and (3.44), we find a constant $k_{1}(\mu,\phi)>0$ such that
(3.54) $\Big{\|}v_{t}^{\varepsilon,2}-\frac{1}{\varepsilon}\mathbb{E}[V(X_{t}^{\mu})(R_{t}^{\varepsilon}-1)]\Big{\|}_{\mathbb{B}}\leq k_{1}(\mu,\phi){{\alpha(\varepsilon)}},\ \ t\in[0,T],\ \varepsilon\in(0,1).$
By (3.14) for $v_{t}=v_{t}(\mu,\phi),\tilde{v}_{t}=\tilde{v}_{t}(\mu,\phi)$, and by the definition of $\hat{v}_{t}$ in (3.35), we have
$\hat{v}_{t}$ $=I_{t}^{V}(\mu,\phi)+\mathbb{E}\bigg{[}V(X_{t}^{\mu})\int_{0}^{t}\Big{\langle}\zeta_{s}(X_{s}^{\mu})\big{\{}N_{s}(\hat{v})+\tilde{N}_{s}(\tilde{v})\big{\}},\ \text{\rm{d}}W_{s}\Big{\rangle}\bigg{]}.$
Combining this with (3.33), (3.38), (3.52), (3.51), (3.54) and the same argument to (3.27), we find a constant $k_{2}(\mu,\phi)$ and a measurable function $\tilde{h}:(0,T]\times[0,1]\rightarrow(0,\infty)$ with
(3.55) $\sup_{\varepsilon\in(0,1],t\in(0,T]}\frac{\sqrt{t}}{\alpha(t^{\frac{1}{2}})}\tilde{h}_{t}(\varepsilon)\leq k_{2}(\mu,\phi),\ \ \lim_{\varepsilon\rightarrow 0}\sup_{t\in(0,T]}\frac{\sqrt{t}}{\alpha(t^{\frac{1}{2}})}\tilde{h}_{t}(\varepsilon)=0,$
such that
$\|v_{t}^{\varepsilon}-\hat{v}_{t}\|_{\mathbb{B}}$
(3.56) $\leq\tilde{h}_{t}(\varepsilon)+\bigg{\|}\mathbb{E}\bigg{[}V(X_{t}^{\mu})\int_{0}^{t}\Big{\langle}\zeta_{s}(X_{s}^{\mu})\big{\{}N_{s}(v^{\varepsilon}-\hat{v})+\tilde{N}_{s}(\tilde{v}^{\varepsilon}-\tilde{v})\big{\}},\ \text{\rm{d}}W_{s}\Big{\rangle}\bigg{]}\bigg{\|}_{\mathbb{B}}$
$\leq\tilde{h}_{t}(\varepsilon)+k_{2}(\mu,\phi)\bigg{(}\int_{0}^{t}\Big{\{}\|v_{s}^{\varepsilon}-\hat{v}_{s}\|_{\mathbb{B}}^{2}+\big{(}\mathbb{E}[|\tilde{v}_{s}^{\varepsilon}-\tilde{v}_{s}|^{k}]\big{)}^{\frac{2}{k}}\Big{\}}\text{\rm{d}}s\bigg{)}^{\frac{1}{2}},\ \ t\in(0,T].$
(f) Proof of (3.35). By (3.49)-(3.3) and Fatou’s lemma, we see that
$\begin{split}&\beta_{t}:=\limsup_{\varepsilon\rightarrow 0}\mathbb{E}\Big{[}\sup_{s\in[0,t]}|\tilde{v}_{s}^{\varepsilon}-\tilde{v}_{s}|^{k}\Big{]},\ \ \tilde{\beta}_{t}:=\limsup_{\varepsilon\rightarrow 0}\sup_{s\in(0,t]}\frac{\sqrt{s}}{\alpha(s^{\frac{1}{2}})}\|v_{s}^{\varepsilon}-\hat{v}_{s}\|_{\mathbb{B}}\end{split}$
satisfy
(3.57) $\begin{split}&\beta_{t}\leq\mathbb{E}\bigg{[}\limsup_{\varepsilon\rightarrow 0}\mathbb{E}\Big{[}\sup_{s\in[0,t]}|\tilde{v}_{s}^{\varepsilon}-\tilde{v}_{s}|^{k}\Big{|}\mathscr{F}_{0}\Big{]}\bigg{]}\\ &\leq c\bigg{(}\int_{0}^{t}\frac{\alpha(s^{\frac{1}{2}})^{2}}{s}\limsup_{\varepsilon\rightarrow 0}\frac{s}{\alpha(s^{\frac{1}{2}})^{2}}\|v_{s}^{\varepsilon}-\hat{v}_{s}\|_{\mathbb{B}}^{2}\text{\rm{d}}s\bigg{)}^{\frac{k}{2}}\\ &\qquad\quad+c\bigg{(}\int_{0}^{t}\Big{\{}\limsup_{\varepsilon\rightarrow 0}\mathbb{E}[|\tilde{v}_{s}^{\varepsilon}-\tilde{v}_{s}|^{k}]\Big{\}}^{2}\text{\rm{d}}s\bigg{)}^{\frac{1}{2}}\\ &\leq c\bigg{(}\int_{0}^{t}\frac{\alpha(s^{\frac{1}{2}})^{2}}{s}\,\tilde{\beta}_{s}^{2}\text{\rm{d}}s\bigg{)}^{\frac{k}{2}}+c\left(\int_{0}^{t}\beta_{s}^{2}\text{\rm{d}}s\right)^{\frac{1}{2}},\ \ t\in[0,T].\end{split}$
Noting that (3.16), (3.34) and (3.45) imply that $\tilde{\beta}_{t}$ satisfies
$\sup_{t\in(0,T]}\tilde{\beta}_{t}\leq\sup_{\varepsilon\in(0,1]}\sup_{s\in(0,T]}\frac{\sqrt{s}}{\alpha(s^{\frac{1}{2}})}\|v_{s}^{\varepsilon}-\hat{v}_{s}\|_{\mathbb{B}}=:\tilde{c}(\mu,\phi)<\infty,$
so that by Fatou’s lemma in (3.3) we derive from (3.55) that
$\tilde{\beta}_{t}^{2}\leq Ck_{2}(\mu,\phi)^{2}\int_{0}^{t}\Big{\{}\frac{\alpha(s^{\frac{1}{2}})^{2}}{s}\tilde{\beta}_{s}^{2}+\beta_{s}^{\frac{2}{k}}\Big{\}}\text{\rm{d}}s,\ \ t\in(0,T],$
where by (2.2),
$C:=\sup_{t\in(0,T]}\frac{t}{\alpha(t^{\frac{1}{2}})^{2}}<\infty.$
Combining this with (3.57), $\int_{0}^{T}\frac{\alpha(t^{\frac{1}{2}})^{2}}{t}\text{\rm{d}}t<\infty$, and applying Gronwall’s inequality, we prove (3.35) and hence finish the proof. ∎
We are now ready to prove Theorem 3.2.
Proof of Theorem 3.2.
By (3.10) and (3.12), it suffices to prove that for any $t\in(0,T]$ and $f\in\mathscr{B}_{b}(\mathbb{R}^{d})$,
(3.58) $\begin{split}\lim_{\varepsilon\downarrow 0}\frac{P_{t}^{\mu_{\varepsilon}}f(\mu_{\varepsilon})-P_{t}^{\mu}f(\mu_{\varepsilon})}{\varepsilon}=\mathbb{E}\bigg{[}f(X_{t}^{\mu})\int_{0}^{t}\Big{\langle}\zeta_{s}(X_{s}^{\mu})\big{\{}N_{s}+\tilde{N}_{s}\big{\}},\ \text{\rm{d}}W_{s}\Big{\rangle}\bigg{]}.\end{split}$
Let $R_{t}^{\varepsilon}$ be in (3.39). By (3.41) for $f$ replacing $V$, we obtain
(3.59) $\frac{P_{t}^{\mu_{\varepsilon}}f(\mu_{\varepsilon})-P_{t}^{\mu}f(\mu_{\varepsilon})}{\varepsilon}=\frac{1}{\varepsilon}\mathbb{E}\big{[}f(X_{t}^{\mu_{\varepsilon},\mu})(R_{t}^{\varepsilon}-1)\big{]},\ \ t\in(0,T].$
Noting that (3.53) implies
$\lim_{\varepsilon\rightarrow 0}\mathbb{E}\Big{[}\sup_{t\in[0,T]}|X_{t}^{\mu_{\varepsilon},\mu}-X_{t}^{\mu}|^{k}\Big{]}=0,$
while $(B_{2})$, Lemma 3.5 and (3.39) lead to
$\lim_{\varepsilon\rightarrow 0}\frac{R_{t}^{\varepsilon}-1}{\varepsilon}=\int_{0}^{t}\Big{\langle}\zeta_{s}(X_{s}^{\mu})\big{\{}N_{s}+\tilde{N}_{s}\big{\}},\ \text{\rm{d}}W_{s}\Big{\rangle}$
in $L^{2}(\mathbb{P})$, by taking $\varepsilon\rightarrow 0$ in (3.59) we deduce (3.58) for $f\in C_{b}(\mathbb{R}^{d})$. By an approximation argument as in [29, Proof of (2.3)] for $f\in\mathscr{B}_{b}(\mathbb{R}^{d})$, this implies (3.59) for $f\in\mathscr{B}_{b}(\mathbb{R}^{d}).$